### v1

# Efficient Multi-agent Navigation with Lightweight DRL Policy

Abstract
In this article, we present an end-to-end collision avoidance policy based on deep reinforcement learning (DRL) for multi-agent systems, demonstrating encouraging outcomes in real-world applications. In particular, our policy calculates the control commands of the agent based on the raw LiDAR observation. In addition, the number of parameters of the proposed basic model is 140,000, and the size of the parameter file is 3.5 MB, which allows the robot to calculate the actions from the CPU alone. We propose a multi-agent training platform based on a physics-based simulator to further bridge the gap between simulation and the real world. The policy is trained on a policy-gradients-based RL algorithm in a dense and messy training environment. A novel reward function is introduced to address the issue of agents choosing suboptimal actions in some common scenarios. Although the data used for training is exclusively from the simulation platform, the policy can be successfully transferred and deployed in real-world robots. Finally, our policy effectively responds to intentional obstructions and avoids collisions. The website is available at https://sites.google.com/view/xingrong2024efficient/%E9%A6%96%E9%A1%B5.
keywords:
Multi-agent systems, distributed collision avoidance, reactive and sensor-based planning, simulation to real gap.
1 Introduction
Multi-agent systems are increasingly used in robotics and artificial intelligence, such as industrial automation (e.g., ), material transportation (e.g., Xie et al. (2024)), and agricultural automation (e.g., Dutta et al. (2021)). The main challenge is implementing an efficient, robust, and safe navigation algorithm for multiple robots so that the robot can safely and efficiently complete point-to-point movement.
The traditional approaches for solving the problem of multi-agent navigation are primarily categorized into two types. One of the approaches is the centralized approach. These approaches calculate trajectories for each agent through a central server using the state of all agents and their respective workspaces (e.g., Tang et al. (2018), Zhou et al. (2021)). However, the computational cost associated with these approaches escalates as the number of agents increases, along with the maintenance of a more extensive communication network. In addition, once the central server is attacked, all agents will stop working. The other approach is the decentralized approach, which is deployed to each agent, empowering them with the capability to make decisions (e.g., Reciprocal Velocity Obstacle (RVO, Fiorini and Shiller (1998)), Optimal Reciprocal Collision Avoidance (ORCA, van den Berg et al. (2011))). However, these algorithms rely on a perfect perception of the environment and the state of the other agents, posing challenges for their application in larger-scale scenarios and deployment onto real-world robots. Numerous variants have been proposed to address these limitations (e.g., Alonso-Mora et al. (2010), Godoy et al. (2016)). Additionally, RVO and its variants often necessitate tuning numerous hyperparameters to achieve optimal performance, and the adjustment of these parameters is highly dependent on the specific scenario.
Learning-based approaches have been proposed for robot navigation in recent years. Some approaches use imitation learning to learn from expert data and work successfully (e.g., Karnan et al. (2022), Yan et al. (2023)). However, the effectiveness of imitation learning depends on expert data quality, and it is challenging to handle scenarios out of the expert data. Other approaches employ reinforcement learning (RL) or deep reinforcement learning (DRL) to train policy. Some RL-based approaches achieve desired outcomes by modifying the reward function (e.g., Xie and Dames (2023), Han et al. (2022)). These approaches can yield great results in corresponding scenarios, but some suffer from overly redundant reward function designs and struggle to handle unforeseen scenarios. Some approaches train agents in specific scenarios to learn more diverse policies (e.g., Fan et al. (2020), Xue and Chen (2024)). Learning in more specific scenarios can enrich policies but demands higher standards for scenario design. Besides, there are efforts to bridge the gap between simulation and the real world by using progressive networks or adding randomness to the simulator (e.g., Rusu et al. (2016), Xie et al. (2023)). Such approaches enable better deployment of policies learned from simulators on real robots, yet they do not fully leverage physics-based simulators. Finally, as the robot in a decentralized system can only perceive local observations during navigation, the decentralized robot navigation can be defined as the partially observable Markov decision process (POMDP). Most policies use convolutional neural networks (CNN) to encode perceptual information (e.g., Liu et al. (2020), Chen et al. (2020)), which results in the loss of historical information during navigation. Some approaches propose using Long Short-Term Memory (LSTM, Hochreiter and Schmidhuber (1997)) to record memory, providing more comprehensive information for navigation (e.g., Liu et al. (2021), Cui et al. (2022)).
In this paper, we propose a lightweight end-to-end distributed multi-agent navigation policy that maps raw LiDAR observation into action commands for the agents. We use the LiDAR sensor in perception for the following reasons (Zhu and Zhang (2021)): 1. the gap of the LiDAR sensor between the simulation and the real world is close; 2. the small measurement dimension of LiDAR is beneficial for simulation-to-real transfer. Furthermore, we introduce a physics-based simulation platform tailored for multi-agent training to bridge the gap between simulation and the real world. We have successfully deployed our policy in simulation and real-world experiments and tested the policy’s performance in single-agent and multi-agent scenarios. Additionally, the policy can avoid intentional obstructions, like a rapid-moving obstacle controlled by a human.
Contribution. Our main contributions are as follows.
  • We propose a multi-agent training platform based on a three-dimensional (3-D) simulator with a physical engine to bridge the gap between simulation and the real world. We compare a two-dimensional (2-D) simulator without a physics engine and a 3-D simulator with a physics engine with real-world environments. The result proves that a 3-D simulator with a physics engine can provide better randomness for agent training.
  • We propose a lightweight end-to-end distributed multi-agent navigation policy, which utilizes the hidden state in the LSTM to alleviate the information loss in the POMDP. Besides, an attention structure based on the lightweight policy is proposed to improve the policy’s performance in multi-agent scenarios. Both policies are trained in a stochastic environment and can be directly deployed to real-world robots.
  • We propose a novel reward function to solve the problem: the agent may choose suboptimal actions in some common scenarios, such as passing through narrow gaps and multi-agent travels in the opposite direction.
The rest of this article is organized as follows. Section 2 is a brief review of the related work. Section 3 introduces the notation and formulation used in this article. Section 4 discusses the policy’s structure, the RL setup for training multi-agent navigation policy, and the simulator platform. Section 5 presents the policy’s performance in the simulator and real-world experiments. Finally, we conclude our work and propose directions for future improvement in Section 6.
2 Related work
In this section, we briefly describe some related works of multi-agent navigation. These works can be categorized into trajectory-based, reaction-based, and learning-based approaches. The learning-based approaches can be further divided into supervised learning and RL. Finally, we discuss the approaches to bridge the gap between simulation and the real world.
2.1 Trajectory-based approaches
Trajectory-based approaches generally require detailed environmental information to plan a feasible trajectory for the robot (e.g., Model Predictive Control (MPC, Keviczky et al. (2004)) and Dynamic Window Approach (DWA, Fox et al. (1997))). In recent years, new theories, such as safe corridors, buffered Voronoi cells (BVC), and Graphs of Convex Sets (GCS), have been introduced to transform the navigation problem into a convex optimization problem. Park et al. (2023) utilizes a dynamic linear safe corridor while introducing mode-based subgoal planning to tackle planning problems in complex environments. They also use collision alert propagation and escape point planning to deal with emergencies. Zhu et al. (2022) constructs a chance-constrained buffered uncertainty-aware Voronoi cell (B-UAVC) for each robot and ensures the safety of robot navigation by specifying thresholds. Marcucci et al. (2023) proposes a framework that enables convex optimization to effectively and reliably plan trajectories around obstacles through a practical convex relaxation programming problem by using GCS. Although the trajectory-based algorithm can obtain the optimal trajectory, it will lead to large computational costs or difficult convex optimization when applied in large or unstructured scenarios.
2.2 Reaction-based approaches
Reaction-based approaches accomplish navigation tasks by continuously sensing and responding to the surrounding environment, enabling them to adapt to dynamically changing conditions. Reciprocal velocity obstacles (RVO) framework (Fiorini and Shiller (1998)) maps the dynamic environment into the robot velocity space and generates safe control velocities from these velocity constraints. The ORCA framework (van den Berg et al. (2011)) follows the principles of RVO while resolving the jitter problem in RVO and improves the obstacle avoidance efficiency. Moreover, Alonso-Mora et al. (2010) further guarantees smooth, collision-free motions under non-holonomic constraints. However, these approaches struggle to be implemented in the real world because they require a perfect perception of the agent’s surroundings. In response to these limitations, some variants have been developed. Godoy et al. (2016) enhances environmental awareness through communication among agents, allowing for more accurate and timely adjustments based on shared information. Arul and Manocha (2021) incorporates Buffered Voronoi Cells to implement a more conservative obstacle avoidance policy, ensuring a safe distance from obstacles and other agents. To some extent, these variants have solved the perception problem. Still, ORCA and its variants are sensitive to parameters, and the corresponding parameters are different in different scenarios, which makes these approaches difficult to adjust in practical applications.
2.3 Supervised learning-based approaches
With the advancement of machine learning, numerous studies utilize it to solve the robot navigation problem. Supervised learning-based approaches train the policy with extensive datasets, either through experts or traditional algorithms, and enable agents to learn how to map observations into actions. Yan et al. (2023) adopts an imitation learning approach, training a mapless navigation policy solely from 2-D LiDAR inputs, incorporating an aggregated dataset with additional safety enhancements. Karnan et al. (2022) utilizes imitation learning to devise a purely vision-based navigation algorithm, successfully learning navigation policies from a single video demonstration collected by physically distinct agents. Cho and Chung (2024) uses a risk-sensitive approach to prioritize learning on scarce but high-risk terrain through self-supervised learning and enriches reliable data for inexperienced terrain with pseudo-labels. Though supervised learning-based approaches have achieved significant breakthroughs in robot navigation, they struggle to cope with scenarios not represented in the training dataset. Furthermore, the reliance on expert knowledge inherently limits the policies from surpassing the boundaries of that preexisting knowledge.
2.4 Reinforcement learning-based approaches
RL can be applied well to navigation tasks since the robot navigation problem is a POMDP problem. RL-based multi-agent navigation approaches learn from the interaction, and the quality of the policy depends on the design of the reward function, the interaction quality, and the neural network’s structure.
Generally, the reward function of the navigation problem is similar to the potential field function. Chen et al. (2017b) proposes a dense reward function that can guide reinforcement learning concisely and effectively. However, the reward function will cause the agent to learn suboptimal actions, which we will discuss in Section 4. Some approaches define the reward function through other rules. Chen et al. (2017a) develops a time-efficient navigation policy that respects public social norms by inducing the right-handed rules in the reward function. Xie and Dames (2023) introduces a velocity obstacle term into the reward function, enabling mobile robots to navigate autonomously in spaces filled with static obstacles and dense pedestrian traffic.
Some approaches train the policy in specific international scenarios. Li et al. (2019) proposes an advanced version of the Social Attention Reinforcement Learning (SARL) algorithm, achieving human-aware navigation in indoor environments. Yao et al. (2020) employs a grid map-based sensor fusion approach to map three frames of egocentric grid maps and the robot’s relative local target positions into low-level robotic control commands. Huang et al. (2024) provides the agent with a curriculum and a replay buffer of clipped collision segments, enhancing performance in obstacle-rich environments.
Regarding the neural network structure, some policies use CNNs to encode perceptual information, such as Liu et al. (2020), Guldenring et al. (2020), Chen et al. (2021). To enhance the policy’s utilization of past states, some approaches harness the hidden state structure of RNNs to store historical information. Liu et al. (2021) presents the Decentralized Structured Recurrent Neural Network (DS-RNN), which uses spatial and temporal relationships for robotic decision-making in swarm navigation. Xue and Chen (2024) introduces a novel multi-agent recurrent deterministic policy gradient algorithm and leverages LSTM to improve action value predictions using historical exploration information.
Some traditional control approaches have also been employed to enhance the safety of RL training. Cheng et al. (2019) utilizes a model-based controller leveraging Control Barrier Functions (CBFs) and online learning of unknown system dynamics to ensure safety during the learning phase. Emam et al. (2022) incorporates safety as a differentiable robust control barrier function layer within a model-based RL framework and proposes a modular approach to learning potential reward-driven tasks independently of safety constraints.
2.5 Simulation-to-real gap
As training is inefficient in the real world and easily damages the robot, RL-based approaches are usually trained in simulators. Kahn et al. (2021) proposed an approach to collect data in reality and train policy that can navigate in real-world urban and off-road environments with geometrically distracting obstacles. However, the gap between simulations and real-world environments remains a crucial barrier preventing the widespread adoption of RL-based approaches in real robots. Several approaches have been proposed to address this issue and bridge the simulation-to-reality gap. Rusu et al. (2016) utilizes progressive networks to bridge the reality gap and transfer policies learned in simulation to the real world. Tobin et al. (2017) investigates domain randomization, a straightforward technique for training models on simulated images that transfers to real images by randomizing renderings within the simulator. García and Molina (2019) proposes an approach to evaluate LiDAR sensors and navigation and obstacle avoidance capabilities in simulated scenarios using a real drone platform. Fan et al. (2020) proposes a hybrid control RL approach, which initially learns an end-to-end general obstacle avoidance policy through RL. Subsequently, hybrid control is incorporated into this policy to improve its feasibility and robustness in real-world environments. This approach leverages the strengths of both RL for learning complex behaviors and hybrid control to ensure reliability and stability in practical applications. Shi et al. (2023) proposed a domain adaptation approach based on a cyclic consistent generative adversarial network and a force control transfer approach to bridge the reality gap.
Our approach further bridges the gap between simulation and the real-world environment by designing a training platform based on a physics-based simulator. We also propose a novel policy network based on LSTM and attention mechanism, which is state of the art (SOTA) of the navigation based on pure LiDAR sensor. Furthermore, a novel reward function is designed to solve some common scenarios, such as agent passes through narrow gaps.
3 Problem formulation
In this section, we will introduce the notation of this work and our reward function.
3.1 Notation of reinforcement learning
This work models the multi-agent navigation process as POMDP. The POMDP can be defined as $<\mathcal{I},\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{R},\mathcal{T},done,\gamma>$, where =$\mathcal{I}=\{1,...,N\}$ denotes $N$ agents, $\mathcal{S}$ denotes the set of states of all agents, $O=\{\mathcal{O}_{1},...,\mathcal{O}_{N}\}$ denotes the set of observable state of agents, $\mathcal{A}$ denotes the set of actions, $\mathcal{R}$ denotes the reward function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ denotes the transfer function and $done$ indicates whether an episode has terminated in time or collided with an obstacle, and $\gamma\in[0,1)$ denotes the discount factor. At each time step $t$, the agent’s observation is $o^{t}=[o_{s}^{t},o_{g}^{t},o_{v}^{t}]$, where $o_{s}^{t}\in\mathbb{R}^{5\times L_{n}}$ is formed by concatenating 5 consecutive frames of LiDAR data, with $L_{n}$ being the measurement dimension of the LiDAR data; $o_{g}^{t}\in\mathbb{R}^{2}$ consists of the distance from the agent’s current position to the goal point, and the angle between the agent’s orientation and the direction towards the goal point; $o_{v}^{t}=[v_{i}^{t},\omega_{i}^{t}]\in\mathbb{R}^{2}$ consists of the agent’s linear and angular velocities. The agent $i$ receives $o_{i}^{t}$ and executes action $a_{i}^{t}$ according to policy $\pi_{\theta}(a^{t}_{i}|o^{t}_{i})$, where $\theta$ is the policy parameter. The policy $\pi_{\theta}(a^{t}_{i}|o^{t}_{i})$ is a Gaussian distribution, with the mean calculated by a policy network. The action $a^{t}_{i}=[v^{t}_{i},\omega^{t}_{i}]$ is sampled from the policy $\pi_{\theta}(a^{t}_{i}|s^{t}_{i})$. The state $s_{t}$ takes the action set ${a_{i}^{t},...,a_{N}^{t}}$ to transfer to next state $s_{t+1}$ according to $\mathcal{T}(s_{t+1}|s_{t},a_{i}^{t},...,a_{N}^{t})$. Also, the next observable state set is updated as $\mathcal{O}=\mathcal{T}(\{o_{1},...o_{N}\}|s_{t},a_{i}^{t},...,a_{N}^{t})$, and the reward set is obtained by =$\mathcal{R}=\{r(s_{1}^{t},a_{1}^{t}),...,r(s_{N}^{t},a_{N}^{t})\}$. A trajectory can be represented as $<o^{1},a^{1},r^{1},o^{2},a^{2},r^{2},...,s^{T},a^{T},r^{T}>$, where $T$ represents the number of interaction steps. Each agent $i$ aims to maximize the discounted future reward: $G_{i}^{t}=\sum_{t=0}^{T}\gamma^{t}r_{i}^{t}$. The state-value function is defined as $V^{\pi}(s)=\mathbb{E}[\sum_{t=0}^{T}\gamma^{t}R^{t}|s^{1}=s]$, and the action-value function is given by $Q^{\pi}(s,a)=\mathbb{E}[R(s,a)+\gamma V^{\pi}(s)|s^{1}=s,a^{1}=a]$. A action is chosen as $a={\rm arg\ max}_{a}Q(s,a)$.
3.2 Reward function
The dense reward function used in RL-based navigation algorithms consists of distance reward, obstacle penalty, and efficiency reward. An example of such a reward function is:
$\begin{array}[]{l}R=r_{g}+r_{c}+r_{other},\\ r_{g}=\left\{\begin{array}[]{ll}r_{{arrival}}\qquad\qquad\quad\ \ \text{if }\left\|p_{i}^{t}-g_{i}\right\|<0.1,\\ w_{g}\left(\left\|p_{i}^{t-1}-g_{i}\right\|-\left\|p_{i}^{t}-g_{i}\right\|\right)\quad\text{otherwise,}\end{array}\right.\\ r_{c}=\left\{\begin{array}[]{ll}r_{{collision}}\qquad\qquad\quad\text{if }d_{i}^{t}<0,\\ w_{c}(0.1-d_{i}^{t}/2)\quad\text{ if }d_{i}^{t}<0.2,\\ 0\qquad\qquad\qquad\quad\text{ otherwise,}\end{array}\right.\\ \end{array}$ (1)
where $r_{other}$ is the other reward that varies in different approaches, $r_{g}$ means the distance reward, which is computed as the difference in the distance between the robot and the goal point at the current time step and the previous time step, $r_{c}$ means the obstacle penalty, $w_{g}$ and $w_{c}$ are the weight of $r_{g}$ and $r_{c}$ respectively, $d_{i}^{t}$ means the smallest distance between agent $i$ and obstacles at time step $t$. However, this reward function setup may make agents learn conservative policy. Additionally, many works impose control constraints, such as velocity and angular velocity, on $r_{other}$ to maintain action coherence. Too many constraints may lead to instability in the training process and difficulty adjusting the weights between various terms. Moreover, in some scenarios, the reward function does not adequately indicate optimal actions.
(a) Agent passes through narrow gaps.
(b) Multi-agent traveling in opposite directions.
Figure 1: Cases that are difficult to handle with the classical reward function.
Figure 2: The weight of the scan penalty of the $r^{\prime}_{c}$.
Figure 3: The overall framework diagram of our approaches is presented. The approaches can be divided into trajectory collection and policy update. During trajectory collection, each agent collects the trajectory independently and stores it in a replay buffer. The simulation platform will return each agent’s reward and observation after an action is executed. After a specific number of iterations, the trajectories stored in the replay buffer will be used to calculate the loss, and the network will then be updated based on the loss through gradient descent. The two stages alternate until a specific number of episodes is reached or the policy converges.
As shown in Figure 1, when an agent needs to navigate through the gap between two obstacles, it may incur penalties due to $r_{c}$ when approaching them, even though traversing the gap is the optimal action. Consequently, the agent may choose to circumnavigate around the obstacles instead. Besides, when multiple agents approach each other, they may only focus on the closest agent or obstacle to itself, neglecting potentially threatening agents that are further away but pose a real danger, which leads to unsafe navigation. To address the issues above, we have introduced modifications to the $r_{c}$ component, specifically by:
$\begin{array}[]{l}r^{\prime}_{c}=\left\{\begin{array}[]{ll}r_{collision}&\text{if }d_{i}^{t}<0,\\ w^{\prime}_{c}\left(l-sl_{i}^{t}\right)+r_{c}&\text{ otherwise,}\end{array}\right.\end{array}$ (2)
where $r_{arrival}=5$, $r_{collision}=-5$, $w_{g}=5$, $w^{\prime}_{c}=-0.25$, $w_{c}=-0.03$, $l$ represents the maximum length of the LiDAR rays, $sl_{i}^{t}$ represents the LiDAR information of agent $i$ at time step $t$, $w^{\prime}_{c}$ is a weight vector that varies with angular velocity, with the weight decreasing in $\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-\omega\Delta t)^{2}}{2\sigma^{2}}}$ fashion from $\omega\Delta t$ towards both sides, where $\sigma=0.2$, $\Delta t$ is the time interval. The schematic diagram of $w^{\prime}_{c}$ is shown in Figure 2.
Using $r^{\prime}_{c}$, the policy pays more attention to the direction of the current angular velocity while maintaining the closest distance to the surrounding obstacles. This allows the policy to better cope with the mentioned situations and allows the agent to pay more attention to the choice of angular velocity.
4 Method
In this section, we will introduce the architecture of the policy network, the simulation platform, and the training process with training scenarios. We utilize the hidden state of LSTM to compensate for the lack of information in the observations and propose an attention structure to improve the policy ability in the multi-agent scenario. Besides, we propose a multi-agent simulation platform to bridge the gap between simulation and the real world. Additionally, the multi-agent training process is based on Proximal Policy Optimization (PPO, Schulman et al. (2017)), with modifications in trajectory collection. Finally, we use two scenarios for training: single-agent and multi-agent.
4.1 Network structure
We note that the agent only requires more biased actions toward the goal point when approaching it. In other situations, the agent merely needs to move in a general direction. As such, we clip the distance between the agent and the goal point to enable the policy to better distinguish between these two situations rather than directly using the distance as an input to the network.
Algorithm 1 PPO For Multi-Agent training
1:procedure PPO(environment, hyperparameters)
2:     Initialize policy network $\pi_{\theta}$, value function network $V_{\phi}$, and set hyper-parameters
3:     for $iteration=1,2,...,$ do
4:         for $t=1,2,...,T_{max}$ do
5:              for robot $i=1,2,...$ do
6:                  Run policy network $\pi_{\theta}$, collecting $\{o_{i}^{t},a_{i}^{t},r_{i}^{t}\}$.
7:                  Run value function network, collecting $\{V_{\phi}(o_{i}^{t})\}$
8:              end for
9:         end for
10:         for robot $i=1,2,...$ do
11:              for $t=T_{max},T_{max-1},...,1$ do
12:                  Calculate TD-error: $\delta_{i}^{t}=r_{i}^{t}+\gamma V_{\phi}(o_{i}^{t+1})-V_{\phi}(o_{i}^{t})$
13:                  Estimate advantages using GAE: $A_{i}^{t}=\sum_{t}^{T_{max}}(\gamma\lambda)^{t}\delta_{i}^{t}$
14:                  Calculate $V_{i}^{t}=A_{i}^{t}+V_{\phi}(o_{i}^{t})$
15:              end for
16:         end for
17:         Merge and shuffle the collected trajectories
18:         $\theta_{old}=\theta,\ \phi_{old}=\phi$
19:         for epoch in Epoch do
20:              $L^{{PPO}}=-\mathbb{E}\left(\sum_{i=1}^{N}\sum_{t=1}^{T_{max}}\text{min}\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{{old}}}(a_{i}^{t}|o_{i}^{t})}A_{i}^{t},\text{clip}\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{{old}}}(a_{i}^{t}|o_{i}^{t})},1-\epsilon,1+\epsilon\right)A_{i}^{t}\right)\right)$
21:              $L^{V}=\text{max}\left(\mathbb{E}\left(\sum_{i=1}^{N}\sum_{t=1}^{T_{max}}\left((V_{\phi}(o^{t}_{i})-V_{i}^{t})^{2},\left(\text{clip}\left(V_{\phi}(o_{i}^{t}),V_{\phi_{old}}(o_{i}^{t})-\epsilon,V_{\phi_{old}}(o_{i}^{t})+\epsilon\right)-V_{i}^{t}\right)^{2}\right)\right)\right)$
22:              $L^{{Entropy}}=H(\pi_{\theta}(a^{t}|o^{t}))$
23:              $L^{{Total}}=L^{{PPO}}+\alpha L^{V}+\beta L^{{Entopy}}$
24:              Update the parameters of $\pi_{\theta}$ and $V_{\phi}$ using Adam Optimizer w.r.t $L^{Total}$
25:         end for
26:     end for
27:end procedure
We first only use LSTM to encode the LiDAR observation. The policy is trained in a static environment and works well in static and 6-agent test environments. However, it performs poorly when the robot’s number increases over 6. Thus, we add an attention module to overcome the issue. As shown in Figure 3, we use a three-layer LSTM module to encode the laser measurements $o_{s}^{t}$, with both hidden sizes = 256, and get $S$. Then, $K$ and $V$ are calculated by two different FC layers by using $S$, in which the output sizes = 256. At the same time, $[o_{g}^{t},\ o_{v}^{t}]$ passes through two FC layers, which output sizes = 256, to obtain $Q$. Then, the encoding of $o_{s}^{t}$, $X$, is obtained by multi-head attention:
$\begin{array}[]{l}\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h})W^{O},\\ \text{where }\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),\\ \text{Attention}(Q,K,V)=\text{Softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V,\end{array}$ (3)
where $d_{k}$ is the size of $K$. The final encoding is obtained by the weighted sum of $X$ and $S$ with an adaptive residual adjustment, $\alpha X+(1-\alpha)S$, where $\alpha$ updates with the policy and the value network. In actor and critic, the final encoding is decoded by two fully connected layer (FC) layers with output size = 128 and output size = 2 to get the mean of the policy $[v^{t},\omega^{t}]$, and another two FC layers with output size = 128 and output size = 1 to get the value network $V_{\phi}(o^{t})$. Note that each LSTM layer is followed by a rectified linear unit (ReLU), and so is the FC layer. Additionally, the last FC layer of the actor and critic modules has no ReLU layers.
4.2 Simulation platform
To bridge the gap between simulation and the real world, we propose a multi-agent training platform based on a physics-based simulator such as Pybullet (Coumans and Bai (2016–2021)) and Gazebo. In most 2-D simulators, the actions of the agent are usually perfectly executed, and their positions are usually calculated by $x^{t+1}=x^{t}+v^{t}\Delta t+\frac{1}{2}a^{t}\Delta t^{2}$. In a physics-based simulator, the agent state is computed by the physics engine, i.e., $x^{t+1}=T(x^{t},v^{t},a^{t},f^{t},...)$, where $T$ is the transfer function of the physics engine, $f^{t}$ is the force acting on the agent. The transfer function $T$ makes the execution of the action randomness. A straightforward example is when the straight action executes in different types of simulators, the results differ. As shown in Figure 4. The 2-D simulator executes the action without bias, while the agent in the physics-based simulator deviates from the straight line, and the deviation is greater in real-world robots. This indicates that training on a physics-based simulator may help the agent learn more realistic policies, thereby facilitating the deployment of the policy on actual robots. To better train the policy, we add Gaussian noise to the agent’s perception to increase randomness further.
(a) 2-D simulator.
(b) 3-D simulator.
(c) Real robot.
Figure 4: Comparison between the simulators and the real world. Agents on all platforms execute commands with $v=1\ m/s,\ \omega=0\ rad/s$. The direct travel distance of the agent in the simulator is 10m, and the distance of the real robot is 5 m.
We propose a local replay mechanism for the trajectory collection of the agent’s collision aiming to speed up the collection of multi-agent interaction trajectories. When the agent collides, the agent will be reset back to the state before the $N$ simulation steps so that the agent can quickly accumulate the collision data and speed up the training, where $N=300$ in this article. However, when the local replay mechanism is deployed in a multi-agent scenario, the reset agent may appear in front of the normal agent, causing the normal agent to be disturbed. Sudden disturbances may make the policy conservative. However, we found that in some cases, agents perform slowing or stopping actions to wait for other agents to leave.
Besides, to make the simulator generalize, we adopt the architecture of Gym (Brockman et al. (2016)), and we have diligently maintained the interfaces intending to provide users with a more convenient experience when utilizing the simulator. The simulator supports the random addition of random or queued agents or obstacles and controls the movement of each agent through a action set $\mathcal{A}$. At the end of a simulator step, the simulator returns a observed state set $\mathcal{O}$. The update frequency of the simulator can be modified to meet the policy training at different control frequencies. All of these functionalities are designed for easy reconfiguration. For example, to accommodate users seeking to evaluate Safe Reinforcement Learning (Safe-RL, Achiam and Amodei (2019)), we have reconfigured our custom-designed simulator to include a constrained version by changing two functions of the origin version.
4.3 Training process
We treat distributed multi-agent navigation as plenty of single agents navigating in an unknown environment, meaning each agent collects the trajectory respectively. As shown in Figure 3, the training process can be divided into trajectory collection and policy update. During the trajectory collection, each agent first observes the LiDAR $o_{s}^{t}$, target point $o_{g}^{t}$, and velocity $o_{v}^{t}$. After that, it utilizes the policy and value network to calculate a value $v^{t}$ and an action mean $[v^{t},\omega^{t}]$. From these, an action $a^{t}$ is sampled based on the policy distribution $\pi_{\theta}(a^{t}|o^{t})$ and executed. The simulator then responds with a reward $r^{t}$, the next state $o^{t+1}$, and a termination signal $done^{t}$. The tuple of $<o^{t},a^{t},r^{t},o^{t+1},done^{t},V_{\phi}(o^{t})>$ is stored in replay buffer. This process is repeated until a specified number of iterations have been completed. Subsequently, the trajectories stored in the relay buffer are used to calculate the total loss $L^{Total}$ for updating the policy and value networks. After updating the networks, the replay buffer will be cleared, and the agents collect trajectories again. The two stages alternate until convergence or episode limit is reached.
Alg 1 shows a more detailed training process. In lines 3-9, each agent collects the trajectories respectively and calculates the values using the value network under the same environment. After specific steps, in lines 10-15, according to the trajectories collected by each agent, the corresponding advantages of each trajectory are calculated through General Advantage Estimation (GAE, Schulman et al. (2015)). Then, we merge and shuffle the trajectories to get training data. In lines 17-24, we calculate the policy loss by $L^{{PPO}}=\text{min}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)}A,\text{clip}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)},1-\epsilon,1+\epsilon)A)$, the value loss by $L^{V}=\max((V_{\theta}-V_{targ})^{2},(\operatorname{clip}(V_{\theta},V_{\theta_{old}}-\varepsilon,V_{\theta_{old}}+\varepsilon)-V_{tart})^{2})$, where $V_{targ}=A+V_{\phi}(s)$ and the entropy loss by $L^{{Entropy}}=H(\pi_{\theta}(a^{t}|o^{t}))$. We get the total loss by weighted summing $L^{{Total}}=-L^{{PPO}}+\alpha L^{V}+\beta L^{{Entopy}}$, where $\alpha$ and $\beta$ are hyperparameters. We use the Adam Optimizer (Kingma and Ba (2014)) to update the policy and value network w.r.t $L^{Total}$. The algorithm proceeds to the next iteration until a specified iteration is reached. We also use some tricks to make the training more stable. The tricks include reward scaling (Gao et al. (2022)), mini-batch updating, advantage normalizing, and learning rate decay.
We use two scenarios during training: single-agent and multi-agent. As we clip the distance input to the network, the size of the scenario does not affect the training. Thus, in the single-agent scenario, the initial point and the goal point of navigation are 10 m apart and the multi-agent scenario’s initial and goal navigation points are 16 m apart. The two scenarios are shown in Figure 5.
(a) Single-agent scenario.
(b) Multi-agent scenario.
[…]
Table 1: Experiment with 1 robot and 15 obstacles.
success rate(%) collision rate(%) trap rate (%) average step
IJRR (Fan et al. (2020)) 91.3 8.6 0.1 1383.23
  Linear policy 67.2 2.4 30.4 550.40
LSTM policy 93.1 6.9 0.0 1187.73
LSTM+Attention 95.9 3.9 0.2 1300.78
  (Safe) LSTM+Attention 97.6 2 0.04 1353.63
Table 2: Experiment with 6 robots and 5 obstacles.
success rate(%) collision rate(%) trap rate (%) average step
NH-ORCA (Alonso-Mora et al. (2010)) 98.4 0.33 1.27 1198.82
IJRR (Fan et al. (2020)) 85.03 14.63 0.33 1617.84
LSTM policy 91.89 8.15 0.0 981.01
LSTM+Attention 94.78 4.68 0.53 1805.94
Figure 6: Multi-agent test without obstacle.
The measurement dimension $L_{n}$ of the LiDAR of the policy is different, and the max length of the LiDAR beam is 4 m. We first train the policy with $L_{n}=36$, which works successfully in the single-agent scenario. In the multi-agent scenario, the $L_{n}$ of LSTM policy and LSTM+Attention are 72. Because, if $L_{n}=36$, the agent may lose information when other agents stay in the internal of the LiDAR beam, with the biggest internal = 0.34 m, while $L_{n}=72$, the biggest internal = 0.17 m is enough for most situations. The field of view (FOV) is set to $\pi$ rad. Additionally, we stack the LiDAR observations of 5 consecutive frames to provide more comprehensive observation information. The agent’s velocity is set within the $[0,1]$ m/s, and the angular velocity is set within the $[-\pi,\pi]$ rad/s. The control frequency used in training is 60 Hz, which can make the reward variation of each action more apparent and maintain efficient control. Finally, the maximum number of simulation steps is set to 2500, sufficient for the agent to finish navigation through the optimal path.
5 Experiments
To evaluate the ability of the policy, we test it in both simulation and the real world compared with NH-ORCA (Alonso-Mora et al. (2010)) and the policy of Fan et al. (2020). In the following section, we refer to the policy of Fan et al. (2020) by CNN-policy, and the measurement dimension of the LiDAR $L_{n}=512$. The LSTM+Attention policy is the proposed policy. We conducted an ablation experiment by removing the attention mechanism from the proposed policy, which is LSTM policy. Additionally, we replaced the LSTM layer with a Linear layer and get the Linear policy, in order to validate the effectiveness of the LSTM layer. In the simulation tests, we use some metrics to evaluate the policy. The metrics consist of the success rate, the collision rate, the trap rate, and the average step. The success rate means the agent’s success rate navigates to the goal in time without collision. The collision rate means the rate at which the agent collides with other objects. The trap rate means the rate at which agent cannot reach the goal in time and without collision. The average step is the average simulation step of the agent to reach the goal successfully. We use the simulation step rather than the cost time of navigation because the cost time depends on the calculation speed of the simulator, which cannot exactly reflect the cost time of navigation. Finally, all simulation tests are performed 1000 times.
5.1 Simulation test
[…]
Figure 7: Real-world multi-agent experiment. The red pattern indicates the state of turtlebot3, the blue pattern indicates the state of a human, the arrow represents the direction of velocity, and the circle represents reaching the goal point (turtlebot3) or remaining stationary (human). At t = 0 s, each turtlebot3 navigates to the goal point. A human appears in the environment at t = 5 s, starting to interfere with turtlebot3. The human first forces a turtlebot3 to change its trajectory and crosses in front of two turtlebot3s at t = 10 s and t = 13 s. Then, the human interferes with another turtlebot3 again until it reaches the goal point for the rest of the time. All turtlebot3s reach the goal points in 31 s.
Figure 8: Real-world single-agent experiment. The turtlebot3 first passes through narrow passages, then crosses through the obstacles, and finally reaches the goal point. The turtlebot3 reaches the goal point in 46 s.
5.2 Real-world test
We use turtlebot3 to test the policy’s performance in the real world. As the max linear velocity and angular velocity of the turtlebot3 are 0.2 m/s and 1 rad/s, we scale the output of our policy to fit the ability of turtlebot3. Besides, the control frequency of turtlebot3 is set to 10 Hz. The observation of the LiDAR given by turtlebot3 with parameters are FOV $=2\pi$ rad, and LiDAR beam number = 360. We only use 72 LiDAR beam with FOV $=0.8\pi$ rad. We align the rest of the policy with the simulation settings. We use the wheel odometry and LiDAR sensor on the turtlebot3 for perception. Wheel odometry calculates the robot’s relative position and angle to the goal point, while the LiDAR sensor is employed to gather LiDAR observation. Although it is possible to let each turtlebot3 calculate its action independently, it is unable to install the corresponding library, like Pytorch Paszke et al. (2017), on the onboard Raspberry Pi. Therefore, we control turtlebot3 through the Robot Operating System (ROS, Quigley et al. (2009)) communication protocol by using a host computer.
In the single-agent scenario, we test the performance of the policy through narrow passages. The turtlebot3 passes and avoids obstacles with unknown positions and shapes to reach the goal point. As shown in Figure 8, the turtlebot3 starts at a room and passes through a passage to get in the room with the goal point in 22 s. Then, the turtlebot3 crosses the obstacles to reach the goal point in 24 s. At last, the turtlebot3 finishes the navigation when approaching the goal point.
In the multi-agent scenario, we test the policy’s performance in avoiding other agents and sudden disturbances. A turtlebot3 needs to go through obstacles and avoid other turtlebot3. Besides, a human who wants to interfere with turtlebot3 is a sudden disturbance. As shown in Figure 7, at first, the human appears in front of a turtlebot3 to force it to change its trajectory, and after that, the human directly crosses through two turtlebot3. Then, the human continues to interfere with other turtlebot3 until it reaches its goal point. All turtlebot3s successfully avoid collision with other turtlebot3s and the human.
These two experiments prove that the policies trained on our proposed simulation platform can be directly deployed on real robots, with only adjustments made to the values of output actions based on the real robot’s capabilities.
6 Conclusion
This paper proposes a lightweight, end-to-end multi-agent navigation policy that translates raw LiDAR observation directly into action commands, eliminating the need to distinguish between dynamic and static obstacles and enabling navigation through narrow gaps. A physics-engine-based simulation platform is introduced for multi-agent training, bridging the gap between simulation and reality. Our method, trained using RL, demonstrates robust performance in simulated and real-world environments, even in malicious obstructions.
Key contributions include the proposal of a distributed navigation policy trained in stochastic environments, the utilization of LSTM to mitigate information loss in POMDP, the development of a multi-agent training platform to facilitate idea validation and safe RL, and the introduction of a new reward function to address suboptimal action selection. The policy’s performance is evaluated in simulation and real-world settings, highlighting its effectiveness and potential for future enhancements.
In the future, we will explore extending the policy to more complex environments, incorporating passing through traffic, or expanding the work to Unmanned aerial vehicles (UAVs). We will further refine the reward function and the policy to enhance performance in edge cases, like U-shaped obstacles.
{acks}
This work is supported by Shenzhen Science and Technology Program under Grant RCBS20221008093305007, Grant 20231115141459001, Young Elite Scientists Sponsorship Program by CAST under Grant 2023QNRC001, and High level of special funds (G03034K003) from Southern University of Science and Technology, Shenzhen, China.

### v3

@@
 Abstract
-In this article, we present an end-to-end collision avoidance policy based on deep reinforcement learning (DRL) for multi-agent systems, demonstrating encouraging outcomes in real-world applications.
-In particular, our policy calculates the control commands of the agent based on the raw LiDAR observation.
+This article presents an end-to-end collision avoidance policy based on deep reinforcement learning (DRL) for multi-agent systems, demonstrating encouraging outcomes in real-world applications.
+In particular, our policy calculates the control commands of the agent based on the raw LiDAR observations.
 In addition, the number of parameters of the proposed basic model is 140,000, and the size of the parameter file is 3.5 MB, which allows the robot to calculate the actions from the CPU alone.
-We propose a multi-agent training platform based on a physics-based simulator to further bridge the gap between simulation and the real world.
+We propose a multi-agent training platform based on a physics-enhanced simulator to further bridge the gap between simulation and the real world.
 The policy is trained on a policy-gradients-based RL algorithm in a dense and messy training environment.
@@
 Finally, our policy effectively responds to intentional obstructions and avoids collisions.
-The website is available at https://sites.google.com/view/xingrong2024efficient/%E9%A6%96%E9%A1%B5.
-keywords:
-Multi-agent systems, distributed collision avoidance, reactive and sensor-based planning, simulation to real gap.
-1 Introduction
-Multi-agent systems are increasingly used in robotics and artificial intelligence, such as industrial automation (e.g., ), material transportation (e.g., Xie et al.
-(2024)), and agricultural automation (e.g., Dutta et al.
-(2021)).
+The website is aviliable in https://sites.google.com/view/xingrong2024efficient/%E9%A6%96%E9%A1%B5
+Note to Practitioners
+We’ve developed a navigation system that enables robots to navigate collision-free in dynamic environments such as warehouses, factories, and hospitals.
+It leverages LiDAR data to make rapid decisions with minimal computational resources, making it highly efficient for small robots.
+This system is trained in physics-enhanced simulations, ensuring its suitability for real-world applications and exceptional performance in handling challenging scenarios like tight spaces and unexpected obstacles.
+Whether you need robots to move efficiently in a bustling warehouse, transport materials in a factory, or assist in a hospital, this approach provides a practical solution for deploying effective navigation without requiring high-end hardware or intricate configurations.
+Index Terms:
+Deep reinforcement learning, distributed collision avoidance, reactive and sensor-based planning, simulation to real gap.
+I INTRODUCTION
+Multi-agent systems are increasingly used in robotics and artificial intelligence, such as industrial automation [1], material transportation [2].
 The main challenge is implementing an efficient, robust, and safe navigation algorithm for multiple robots so that the robot can safely and efficiently complete point-to-point movement.
@@
 One of the approaches is the centralized approach.
-These approaches calculate trajectories for each agent through a central server using the state of all agents and their respective workspaces (e.g., Tang et al.
-(2018), Zhou et al.
-(2021)).
-However, the computational cost associated with these approaches escalates as the number of agents increases, along with the maintenance of a more extensive communication network.
+These approaches calculate trajectories for each agent through a central server using the state of all agents and their respective workspaces [3].
+However, the computational cost and complexity of these approaches increase significantly with the number of agents, and the maintenance of a more extensive communication network.
 In addition, once the central server is attacked, all agents will stop working.
-The other approach is the decentralized approach, which is deployed to each agent, empowering them with the capability to make decisions (e.g., Reciprocal Velocity Obstacle (RVO, Fiorini and Shiller (1998)), Optimal Reciprocal Collision Avoidance (ORCA, van den Berg et al.
-(2011))).
-However, these algorithms rely on a perfect perception of the environment and the state of the other agents, posing challenges for their application in larger-scale scenarios and deployment onto real-world robots.
-Numerous variants have been proposed to address these limitations (e.g., Alonso-Mora et al.
-(2010), Godoy et al.
-(2016)).
+The other approach is the decentralized approach, which is deployed to each agent, empowering them with the capability to make decisions, such as Reciprocal Velocity Obstacle (RVO) [4] and Optimal Reciprocal Collision Avoidance (ORCA) [5].
+However, these algorithms rely on an idealized perception of the environment and the states of the other agents, posing challenges for their application in larger-scale scenarios and deployment onto real-world robots.
+Numerous variants have been proposed to address these limitations [6], [7].
 Additionally, RVO and its variants often necessitate tuning numerous hyperparameters to achieve optimal performance, and the adjustment of these parameters is highly dependent on the specific scenario.
-Learning-based approaches have been proposed for robot navigation in recent years.
-Some approaches use imitation learning to learn from expert data and work successfully (e.g., Karnan et al.
-(2022), Yan et al.
-(2023)).
+Learning-based approaches have been proposed for robot motion planning in recent years [8].
+Some approaches use imitation learning to learn from expert data and work successfully [9].
 However, the effectiveness of imitation learning depends on expert data quality, and it is challenging to handle scenarios out of the expert data.
 Other approaches employ reinforcement learning (RL) or deep reinforcement learning (DRL) to train policy.
-Some RL-based approaches achieve desired outcomes by modifying the reward function (e.g., Xie and Dames (2023), Han et al.
-(2022)).
-These approaches can yield great results in corresponding scenarios, but some suffer from overly redundant reward function designs and struggle to handle unforeseen scenarios.
-Some approaches train agents in specific scenarios to learn more diverse policies (e.g., Fan et al.
-(2020), Xue and Chen (2024)).
+Some RL-based approaches achieve desired outcomes by modifying the reward function [10], [11].
+These approaches can yield excellent results in corresponding scenarios, but some suffer from redundant reward function designs and struggle to handle unforeseen scenarios.
+Some approaches train agents in specific scenarios to learn more diverse policies [12], [13].
 Learning in more specific scenarios can enrich policies but demands higher standards for scenario design.
-Besides, there are efforts to bridge the gap between simulation and the real world by using progressive networks or adding randomness to the simulator (e.g., Rusu et al.
-(2016), Xie et al.
-(2023)).
-Such approaches enable better deployment of policies learned from simulators on real robots, yet they do not fully leverage physics-based simulators.
+Besides, there are efforts to bridge the gap between simulation and the real world by using progressive networks or adding randomness to the simulator [14], [15].
+Such approaches enable better deployment of policies learned from simulators on real robots, yet they do not fully leverage physics-enhanced simulators.
 Finally, as the robot in a decentralized system can only perceive local observations during navigation, the decentralized robot navigation can be defined as the partially observable Markov decision process (POMDP).
-Most policies use convolutional neural networks (CNN) to encode perceptual information (e.g., Liu et al.
-(2020), Chen et al.
-(2020)), which results in the loss of historical information during navigation.
-Some approaches propose using Long Short-Term Memory (LSTM, Hochreiter and Schmidhuber (1997)) to record memory, providing more comprehensive information for navigation (e.g., Liu et al.
-(2021), Cui et al.
-(2022)).
-In this paper, we propose a lightweight end-to-end distributed multi-agent navigation policy that maps raw LiDAR observation into action commands for the agents.
-We use the LiDAR sensor in perception for the following reasons (Zhu and Zhang (2021)): 1.
-the gap of the LiDAR sensor between the simulation and the real world is close; 2.
-the small measurement dimension of LiDAR is beneficial for simulation-to-real transfer.
-Furthermore, we introduce a physics-based simulation platform tailored for multi-agent training to bridge the gap between simulation and the real world.
-We have successfully deployed our policy in simulation and real-world experiments and tested the policy’s performance in single-agent and multi-agent scenarios.
-Additionally, the policy can avoid intentional obstructions, like a rapid-moving obstacle controlled by a human.
+Most policies use convolutional neural networks (CNN) to encode perceptual information [16], [17], which results in the loss of historical information during navigation.
+Some approaches propose using long short-term memory (LSTM) [18] to record memory, providing more comprehensive information for navigation [19], [20].
+This paper presents a lightweight, end-to-end distributed multi-agent navigation framework that directly maps raw LiDAR observations to agent action commands.
+LiDAR is selected as the primary perception modality due to its minimal simulation-to-reality gap and compact measurement dimensions, which facilitate effective simulation-to-real transfer [21].
+We develop a physics-enhanced simulation platform replicating real-world dynamics to bridge the reality gap in multi-agent training.
+Extensive experimental validation demonstrates the policy’s robustness in simulated and physical environments, including single-agent and multi-agent scenarios.
+Notably, the framework successfully handles adversarial conditions, such as rapidly moving human-controlled obstacles.
 Contribution. Our main contributions are as follows.
-  • We propose a multi-agent training platform based on a three-dimensional (3-D) simulator with a physical engine to bridge the gap between simulation and the real world.
-We compare a two-dimensional (2-D) simulator without a physics engine and a 3-D simulator with a physics engine with real-world environments.
-The result proves that a 3-D simulator with a physics engine can provide better randomness for agent training.
-  • We propose a lightweight end-to-end distributed multi-agent navigation policy, which utilizes the hidden state in the LSTM to alleviate the information loss in the POMDP.
-Besides, an attention structure based on the lightweight policy is proposed to improve the policy’s performance in multi-agent scenarios.
-Both policies are trained in a stochastic environment and can be directly deployed to real-world robots.
-  • We propose a novel reward function to solve the problem: the agent may choose suboptimal actions in some common scenarios, such as passing through narrow gaps and multi-agent travels in the opposite direction.
+  • We develop a physics-enhanced simulation platform that bridges the simulation-to-reality gap.
+Comparisons show the physics-integrated system improves environmental randomness and training effectiveness.
+  • We propose a lightweight distributed framework using LSTM hidden states to address POMDP information loss.
+An integrated attention mechanism enhances coordination, with policies trained in stochastic environments and transferable to real-world robots.
+  • We introduce a reward function addressing classical formulation limitations, resolving suboptimal behaviors in narrow passages and bidirectional multi-agent navigation.
 The rest of this article is organized as follows.
-Section 2 is a brief review of the related work.
-Section 3 introduces the notation and formulation used in this article.
-Section 4 discusses the policy’s structure, the RL setup for training multi-agent navigation policy, and the simulator platform.
-Section 5 presents the policy’s performance in the simulator and real-world experiments.
-Finally, we conclude our work and propose directions for future improvement in Section 6.
-2 Related work
+Section II is a brief review of the related work.
+Section III introduces the notation and formulation used in this article, the proposed reward function, the policy’s structure, the overall framework of the approach, and the physics-enhanced simulator platform.
+Section IV presents the policy’s performance in the simulation and real-world experiments.
+Finally, we conclude our work and propose directions for future improvement in Section V.
+II Related Work
 In this section, we briefly describe some related works of multi-agent navigation.
 These works can be categorized into trajectory-based, reaction-based, and learning-based approaches.
-The learning-based approaches can be further divided into supervised learning and RL.
 Finally, we discuss the approaches to bridge the gap between simulation and the real world.
-2.1 Trajectory-based approaches
-Trajectory-based approaches generally require detailed environmental information to plan a feasible trajectory for the robot (e.g., Model Predictive Control (MPC, Keviczky et al.
-(2004)) and Dynamic Window Approach (DWA, Fox et al.
-(1997))).
+II-A Trajectory-based Approaches
+Trajectory-based approaches generally require detailed environmental information to plan a feasible trajectory for the robot [22].
 In recent years, new theories, such as safe corridors, buffered Voronoi cells (BVC), and Graphs of Convex Sets (GCS), have been introduced to transform the navigation problem into a convex optimization problem.
-Park et al.
-(2023) utilizes a dynamic linear safe corridor while introducing mode-based subgoal planning to tackle planning problems in complex environments.
-They also use collision alert propagation and escape point planning to deal with emergencies.
-Zhu et al.
-(2022) constructs a chance-constrained buffered uncertainty-aware Voronoi cell (B-UAVC) for each robot and ensures the safety of robot navigation by specifying thresholds.
-Marcucci et al.
-(2023) proposes a framework that enables convex optimization to effectively and reliably plan trajectories around obstacles through a practical convex relaxation programming problem by using GCS.
-Although the trajectory-based algorithm can obtain the optimal trajectory, it will lead to large computational costs or difficult convex optimization when applied in large or unstructured scenarios.
-2.2 Reaction-based approaches
+Park [23] utilizes a dynamic linear safe corridor while introducing mode-based subgoal planning to tackle planning problems in complex environments.
+Zhu [24] constructs a chance-constrained buffered uncertainty-aware Voronoi cell (B-UAVC) for each robot and ensures the safety of robot navigation by specifying thresholds.
+Marcucci [25] proposes a framework that enables convex optimization to effectively and reliably plan trajectories around obstacles through a practical convex relaxation programming problem by using GCS.
+Although the trajectory-based algorithm can obtain the optimal trajectory, it will lead to significant computational costs or conplex convex optimization when applied in large or unstructured scenarios.
+II-B Reaction-based Approaches
 Reaction-based approaches accomplish navigation tasks by continuously sensing and responding to the surrounding environment, enabling them to adapt to dynamically changing conditions.
-Reciprocal velocity obstacles (RVO) framework (Fiorini and Shiller (1998)) maps the dynamic environment into the robot velocity space and generates safe control velocities from these velocity constraints.
-The ORCA framework (van den Berg et al.
-(2011)) follows the principles of RVO while resolving the jitter problem in RVO and improves the obstacle avoidance efficiency.
-Moreover, Alonso-Mora et al.
-(2010) further guarantees smooth, collision-free motions under non-holonomic constraints.
+Reciprocal velocity obstacles (RVO) [4] maps the dynamic environment into the robot velocity space and generates safe control velocities from these velocity constraints.
+The ORCA [5] follows the principles of RVO while resolving the jitter problem in RVO and improves obstacle avoidance efficiency.
+Moreover, ORCA-NH [6] further guarantees smooth, collision-free motions under non-holonomic constraints.
 However, these approaches struggle to be implemented in the real world because they require a perfect perception of the agent’s surroundings.
-In response to these limitations, some variants have been developed.
-Godoy et al.
-(2016) enhances environmental awareness through communication among agents, allowing for more accurate and timely adjustments based on shared information.
-Arul and Manocha (2021) incorporates Buffered Voronoi Cells to implement a more conservative obstacle avoidance policy, ensuring a safe distance from obstacles and other agents.
-To some extent, these variants have solved the perception problem.
-Still, ORCA and its variants are sensitive to parameters, and the corresponding parameters are different in different scenarios, which makes these approaches difficult to adjust in practical applications.
-2.3 Supervised learning-based approaches
-With the advancement of machine learning, numerous studies utilize it to solve the robot navigation problem.
-Supervised learning-based approaches train the policy with extensive datasets, either through experts or traditional algorithms, and enable agents to learn how to map observations into actions.
-Yan et al.
-(2023) adopts an imitation learning approach, training a mapless navigation policy solely from 2-D LiDAR inputs, incorporating an aggregated dataset with additional safety enhancements.
-Karnan et al.
-(2022) utilizes imitation learning to devise a purely vision-based navigation algorithm, successfully learning navigation policies from a single video demonstration collected by physically distinct agents.
-Cho and Chung (2024) uses a risk-sensitive approach to prioritize learning on scarce but high-risk terrain through self-supervised learning and enriches reliable data for inexperienced terrain with pseudo-labels.
-Though supervised learning-based approaches have achieved significant breakthroughs in robot navigation, they struggle to cope with scenarios not represented in the training dataset.
-Furthermore, the reliance on expert knowledge inherently limits the policies from surpassing the boundaries of that preexisting knowledge.
-2.4 Reinforcement learning-based approaches
+Still, ORCA and its variants are sensitive to parameters, and the corresponding parameters are different in different scenarios, which makes these approaches challenging to adjust in practical applications.
+II-C Reinforcement Learning-based Approaches
 RL can be applied well to navigation tasks since the robot navigation problem is a POMDP problem.
@@
 Generally, the reward function of the navigation problem is similar to the potential field function.
-Chen et al.
-(2017b) proposes a dense reward function that can guide reinforcement learning concisely and effectively.
-However, the reward function will cause the agent to learn suboptimal actions, which we will discuss in Section 4.
+Chen [26] proposes a dense reward function that can guide reinforcement learning concisely and effectively.
 Some approaches define the reward function through other rules.
-Chen et al.
-(2017a) develops a time-efficient navigation policy that respects public social norms by inducing the right-handed rules in the reward function.
-Xie and Dames (2023) introduces a velocity obstacle term into the reward function, enabling mobile robots to navigate autonomously in spaces filled with static obstacles and dense pedestrian traffic.
+Xie [10] introduces a velocity obstacle term into the reward function, enabling mobile robots to navigate autonomously in spaces filled with static obstacles and dense pedestrian traffic.
 Some approaches train the policy in specific international scenarios.
-Li et al.
-(2019) proposes an advanced version of the Social Attention Reinforcement Learning (SARL) algorithm, achieving human-aware navigation in indoor environments.
-Yao et al.
-(2020) employs a grid map-based sensor fusion approach to map three frames of egocentric grid maps and the robot’s relative local target positions into low-level robotic control commands.
-Huang et al.
-(2024) provides the agent with a curriculum and a replay buffer of clipped collision segments, enhancing performance in obstacle-rich environments.
-Regarding the neural network structure, some policies use CNNs to encode perceptual information, such as Liu et al.
-(2020), Guldenring et al.
-(2020), Chen et al.
-(2021).
-To enhance the policy’s utilization of past states, some approaches harness the hidden state structure of RNNs to store historical information.
-Liu et al.
-(2021) presents the Decentralized Structured Recurrent Neural Network (DS-RNN), which uses spatial and temporal relationships for robotic decision-making in swarm navigation.
-Xue and Chen (2024) introduces a novel multi-agent recurrent deterministic policy gradient algorithm and leverages LSTM to improve action value predictions using historical exploration information.
+Regarding the neural network structure, some policies use CNNs to encode perceptual information [16].
+To improve the policy’s ability to leverage historical states, some approaches utilize the hidden state architecture of RNNs to encode and store past information.
+Liu [19] presents the Decentralized Structured Recurrent Neural Network (DS-RNN), which uses spatial and temporal relationships for robotic decision-making in swarm navigation.
+Xue [13] introduces a novel multi-agent recurrent deterministic policy gradient algorithm and leverages LSTM to improve action value predictions using historical exploration information.
 Some traditional control approaches have also been employed to enhance the safety of RL training.
-Cheng et al.
-(2019) utilizes a model-based controller leveraging Control Barrier Functions (CBFs) and online learning of unknown system dynamics to ensure safety during the learning phase.
-Emam et al.
-(2022) incorporates safety as a differentiable robust control barrier function layer within a model-based RL framework and proposes a modular approach to learning potential reward-driven tasks independently of safety constraints.
-2.5 Simulation-to-real gap
+Cheng [27] utilizes a model-based controller leveraging Control Barrier Functions (CBFs) and online learning of unknown system dynamics to ensure safety during the learning phase.
+II-D Simulation-to-Real Gap
 As training is inefficient in the real world and easily damages the robot, RL-based approaches are usually trained in simulators.
-Kahn et al.
-(2021) proposed an approach to collect data in reality and train policy that can navigate in real-world urban and off-road environments with geometrically distracting obstacles.
 However, the gap between simulations and real-world environments remains a crucial barrier preventing the widespread adoption of RL-based approaches in real robots.
-Several approaches have been proposed to address this issue and bridge the simulation-to-reality gap.
-Rusu et al.
-(2016) utilizes progressive networks to bridge the reality gap and transfer policies learned in simulation to the real world.
-Tobin et al.
-(2017) investigates domain randomization, a straightforward technique for training models on simulated images that transfers to real images by randomizing renderings within the simulator.
-García and Molina (2019) proposes an approach to evaluate LiDAR sensors and navigation and obstacle avoidance capabilities in simulated scenarios using a real drone platform.
-Fan et al.
-(2020) proposes a hybrid control RL approach, which initially learns an end-to-end general obstacle avoidance policy through RL.
+Several methods have been proposed to address this issue and bridge the simulation-to-reality gap.
+Carcia [28] proposes an approach to evaluate Lidar sensors and navigation and obstacle avoidance capabilities in simulated scenarios using a real drone platform.
+Fan [12] proposes a hybrid control RL approach, which initially learns an end-to-end general obstacle avoidance policy through RL.
 Subsequently, hybrid control is incorporated into this policy to improve its feasibility and robustness in real-world environments.
-This approach leverages the strengths of both RL for learning complex behaviors and hybrid control to ensure reliability and stability in practical applications.
-Shi et al.
-(2023) proposed a domain adaptation approach based on a cyclic consistent generative adversarial network and a force control transfer approach to bridge the reality gap.
 Our approach further bridges the gap between simulation and the real-world environment by designing a training platform based on a physics-based simulator.
-We also propose a novel policy network based on LSTM and attention mechanism, which is state of the art (SOTA) of the navigation based on pure LiDAR sensor.
-Furthermore, a novel reward function is designed to solve some common scenarios, such as agent passes through narrow gaps.
-3 Problem formulation
-In this section, we will introduce the notation of this work and our reward function.
-3.1 Notation of reinforcement learning
+We propose a novel policy network based on LSTM and attention mechanism, a state-of-the-art (SOTA) navigation based on a pure LiDAR sensor.
+Furthermore, a novel reward function is designed to solve common scenarios, such as an agent passing through narrow gaps.
+III Method
+In this section, we will introduce the notation of this work, the proposed reward function, the architecture of the policy network, the training process, and the simulation platform with training scenarios.
+III-A Notations and Problem Statement
 This work models the multi-agent navigation process as POMDP.
-The POMDP can be defined as $<\mathcal{I},\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{R},\mathcal{T},done,\gamma>$, where =$\mathcal{I}=\{1,...,N\}$ denotes $N$ agents, $\mathcal{S}$ denotes the set of states of all agents, $O=\{\mathcal{O}_{1},...,\mathcal{O}_{N}\}$ denotes the set of observable state of agents, $\mathcal{A}$ denotes the set of actions, $\mathcal{R}$ denotes the reward function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ denotes the transfer function and $done$ indicates whether an episode has terminated in time or collided with an obstacle, and $\gamma\in[0,1)$ denotes the discount factor.
+The POMDP can be defined as $<\mathcal{I},\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{R},\mathcal{T},done,\gamma>$, where =$\mathcal{I}=\{1,...,N\}$ denotes $N$ agents, $\mathcal{S}$ denotes the set of states of all agents, $\mathcal{O}=\{\mathcal{O}_{1},...,\mathcal{O}_{N}\}$ denotes the set of observable state of agents, $\mathcal{A}$ denotes the set of actions, $\mathcal{R}$ denotes the reward function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ denotes the transfer function and $done$ indicates whether an episode has terminated in time or collided with an obstacle, and $\gamma\in[0,1)$ denotes the discount factor.
 At each time step $t$, the agent’s observation is $o^{t}=[o_{s}^{t},o_{g}^{t},o_{v}^{t}]$, where $o_{s}^{t}\in\mathbb{R}^{5\times L_{n}}$ is formed by concatenating 5 consecutive frames of LiDAR data, with $L_{n}$ being the measurement dimension of the LiDAR data; $o_{g}^{t}\in\mathbb{R}^{2}$ consists of the distance from the agent’s current position to the goal point, and the angle between the agent’s orientation and the direction towards the goal point; $o_{v}^{t}=[v_{i}^{t},\omega_{i}^{t}]\in\mathbb{R}^{2}$ consists of the agent’s linear and angular velocities.
@@
 The action $a^{t}_{i}=[v^{t}_{i},\omega^{t}_{i}]$ is sampled from the policy $\pi_{\theta}(a^{t}_{i}|s^{t}_{i})$.
-The state $s_{t}$ takes the action set ${a_{i}^{t},...,a_{N}^{t}}$ to transfer to next state $s_{t+1}$ according to $\mathcal{T}(s_{t+1}|s_{t},a_{i}^{t},...,a_{N}^{t})$.
-Also, the next observable state set is updated as $\mathcal{O}=\mathcal{T}(\{o_{1},...o_{N}\}|s_{t},a_{i}^{t},...,a_{N}^{t})$, and the reward set is obtained by =$\mathcal{R}=\{r(s_{1}^{t},a_{1}^{t}),...,r(s_{N}^{t},a_{N}^{t})\}$.
-A trajectory can be represented as $<o^{1},a^{1},r^{1},o^{2},a^{2},r^{2},...,s^{T},a^{T},r^{T}>$, where $T$ represents the number of interaction steps.
-Each agent $i$ aims to maximize the discounted future reward: $G_{i}^{t}=\sum_{t=0}^{T}\gamma^{t}r_{i}^{t}$.
-The state-value function is defined as $V^{\pi}(s)=\mathbb{E}[\sum_{t=0}^{T}\gamma^{t}R^{t}|s^{1}=s]$, and the action-value function is given by $Q^{\pi}(s,a)=\mathbb{E}[R(s,a)+\gamma V^{\pi}(s)|s^{1}=s,a^{1}=a]$.
-A action is chosen as $a={\rm arg\ max}_{a}Q(s,a)$.
-3.2 Reward function
+The state $s_{t}$ takes the action set ${a_{i}^{t},...,a_{N}^{t}}$ to transfer to next state $s_{i}^{t+1}$ according to $\mathcal{T}(\mathcal{S}^{t+1}|\mathcal{S}^{t},a_{i}^{t},...,a_{N}^{t})$.
+Also, the next observable state set is updated as $\mathcal{T}(O^{t+1}|\mathcal{O}^{t},a_{i}^{t},...,a_{N}^{t})$, and the reward set is obtained by =$\mathcal{R}=\{r(s_{1}^{t},a_{1}^{t}),...,r(s_{N}^{t},a_{N}^{t})\}$.
+A trajectory of agent $i$ can be represented as $<o_{i}^{1},a_{i}^{1},r_{i}^{1},o_{i}^{2},a_{i}^{2},r_{i}^{2},...,s_{i}^{T},a_{i}^{T},r_{i}^{T}>$, where $T$ represents the number of interaction steps.
+Each agent $i$ aims to maximize the discounted future reward: $G_{i}^{T}=\sum_{t=0}^{T}\gamma^{t}r(s_{i}^{t},a_{i}^{t})$.
+The state-value function is defined as $V^{\pi}(s_{i}^{T})=\mathbb{E}[G_{i}^{T}|s^{1}_{i}]$, and the action-value function is given by $Q^{\pi}(s_{i}^{T},a_{i}^{T})=\mathbb{E}[G_{i}^{T}+\gamma V^{\pi}(s_{i}^{T})|s^{1}_{i},a^{1}_{i}]$.
+An action is chosen as $a={\rm arg\ max}_{a}Q(s,a)$.
+III-B Simulation Platform
+To bridge the gap between simulation and the real world, we propose a multi-agent training platform based on a physics-enhanced simulator such as Pybullet [29] and Gazebo.
+In most 2-D simulators, the actions of the agent are usually perfectly executed, and their positions are usually calculated by an equation like $x^{t+1}=x^{t}+v^{t}\Delta t+\frac{1}{2}a^{t}\Delta t^{2}$.
+In a physics-enhanced simulator, the agent state is computed by the physics engine, i.e., $x^{t+1}=\mathcal{T}(x^{t},v^{t},a^{t},f^{t},...)$, where $\mathcal{T}$ is the transfer function of the physics engine, $f^{t}$ is the force acting on the agent.
+The transfer function $\mathcal{T}$ makes the execution of the action randomness.
+A straightforward example is when the straight action executes in different types of simulators, the results differ.
+the 2-D simulator executes actions without deviation, whereas the agent in the physics-enhanced simulator diverges from the straight path, with even more pronounced deviations observed in real-world robots.
+This suggests that training in a physics-enhanced simulator enables the agent to learn more realistic policies, thereby improving its adaptability and facilitating seamless deployment on actual robots.
+To better train the policy, we add Gaussian noise to the agent’s perception to increase randomness further.
+(a) 2-D simulator
+(b) 3-D simulator
+(c) Real robot
+Figure 1: Comparison between the simulators and the real world.
+Agents on all platforms execute commands with $v=1$ m/s, $\omega=0$ rad/s.
+The direct travel distance of the agent in the simulator is 10 m, and the distance of the real robot is 5 m.
+We propose a local replay mechanism for the trajectory collection of the agent’s collision, aiming to speed up the collection of multi-agent interaction trajectories.
+When the agent collides, the agent will be reset back to the state before the $N$ simulation steps so that the agent can quickly accumulate the collision data and speed up the training, where $N=300$ in this article.
+We propose this mechanism to address: (1) Spatiotemporal Sparsity: collisions in dense multi-agent scenarios are rare, making global reset methods inefficient; (2) Emergent Obstacle Prediction: sudden disturbances induce conservative behaviors (e.g., slowing/stopping), revealing implicit dynamic obstacle modeling; and (3) Asymmetric Adaptation: resetting single agents forces rapid adaptation to dynamic obstacles, simulating real-world pedestrian unpredictability and enhancing generalization.
+Besides, to make the simulator generalize, we adopt the architecture of Gym [30], and we have diligently maintained the interfaces intending to provide users with a more convenient experience when utilizing the simulator.
+The simulator supports the random addition of random or queued agents or obstacles and controls the movement of each agent through an action set $\mathcal{A}$.
+At the end of a simulator step, the simulator returns an observed state set $\mathcal{O}$.
+The update frequency of the simulator can be modified to meet the policy training at different control frequencies.
+All of these functionalities are designed for easy reconfiguration.
+For example, to accommodate users seeking to evaluate Safe Reinforcement Learning (Safe-RL) [31], we have reconfigured our custom-designed simulator to include a constrained version by changing two functions of the origin version.
+Figure 2: The overall framework diagram of our approaches is presented.
+The approaches can be divided into trajectory collection and policy update.
+During trajectory collection, each agent collects the trajectory independently and stores it in a replay buffer.
+After an action is executed, the simulation platform will return each agent’s reward and observation.
+After a specific number of iterations, the trajectories stored in the replay buffer will be used to calculate the loss, and the optimizer will then update the network based on the loss through gradient descent.
+The two stages alternate until a specific number of episodes is reached or the policy converges.
+Algorithm 1 PPO For Multi-Agent Training
+1:Initialize policy network $\pi_{\theta}$, value function network $V_{\phi}$, and set hyper-parameters
+2:for $iteration=1,2,\ldots$ do
+3:     for $t=1,2,\ldots,T_{\text{max}}$ do
+4:         for each robot $i=1,2,\ldots,N$ do
+5:              Run policy network $\pi_{\theta}$ to collect $\{o_{i}^{t},a_{i}^{t},r_{i}^{t}\}$
+6:              Run value function network to collect $\{V_{\phi}(o_{i}^{t})\}$
+7:         end for
+8:     end for
+9:     for each robot $i=1,2,\ldots,N$ do
+10:         for $t=T_{\text{max}},T_{\text{max}}-1,\ldots,1$ do
+11:              Calculate TD-error: $\delta_{i}^{t}=r_{i}^{t}+\gamma V_{\phi}(o_{i}^{t+1})-V_{\phi}(o_{i}^{t})$
+12:              Estimate advantages using GAE: $A_{i}^{t}=\sum_{k=t}^{T_{\text{max}}}(\gamma\lambda)^{k-t}\delta_{i}^{k}$
+13:              Calculate $V_{i}^{t}=A_{i}^{t}+V_{\phi}(o_{i}^{t})$
+14:         end for
+15:     end for
+16:     Merge and shuffle the collected trajectories
+17:     $\theta_{\text{old}}=\theta$, $\phi_{\text{old}}=\phi$
+18:     for each epoch in Epoch do
+19:         Calculate policy loss:
+$L^{P}=-\mathbb{E}\left[\sum_{i=1}^{N}\sum_{t=1}^{T_{\text{max}}}\min\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{\text{old}}}(a_{i}^{t}|o_{i}^{t})}A_{i}^{t},\text{clip}\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{\text{old}}}(a_{i}^{t}|o_{i}^{t})},1-\epsilon,1+\epsilon\right)A_{i}^{t}\right)\right]$
+20:         Calculate value loss:
+$L^{V}=\max\left(\mathbb{E}\left[\sum_{i=1}^{N}\sum_{t=1}^{T_{\text{max}}}\left((V_{\phi}(o_{i}^{t})-V_{i}^{t})^{2},\left(\text{clip}(V_{\phi}(o_{i}^{t}),V_{\phi_{\text{old}}}(o_{i}^{t})-\epsilon,V_{\phi_{\text{old}}}(o_{i}^{t})+\epsilon)-V_{i}^{t}\right)^{2}\right)\right]\right)$
+21:         Calculate entropy loss: $L^{\text{Entropy}}=\mathbb{E}[H(\pi_{\theta}(a_{i}^{t}|o_{i}^{t}))]$
+22:         Calculate total loss: $L^{\text{Total}}=L^{P}+\alpha L^{V}+\beta L^{\text{Entropy}}$
+23:         Update the parameters of $\pi_{\theta}$ and $V_{\phi}$ using Adam Optimizer with respect to $L^{\text{Total}}$
+24:     end for
+25:end for
+III-C Training Process
+We treat distributed multi-agent navigation as plenty of single agents navigating in an unknown environment, meaning each agent collects the trajectory respectively.
+As shown in Fig.
+2, the training process can be divided into trajectory collection and policy update.
+During the trajectory collection, each agent first observes the LiDAR $o_{s}^{t}$, target point $o_{g}^{t}$, and velocity $o_{v}^{t}$.
+After that, it utilizes the policy and value network to calculate an action mean $[v^{t},\omega^{t}]$ and a value $v^{t}$.
+From these, an action $a^{t}$ is sampled based on the policy distribution $\pi_{\theta}(a^{t}|o^{t})$ and executed.
+The simulator then responds with a reward $r^{t}$, the next observable state $o^{t+1}$, and a termination signal $done^{t}$.
+The tuple of $<o^{t},a^{t},r^{t},o^{t+1},done^{t},V_{\phi}(o^{t})>$ is stored in replay buffer.
+This process is repeated until a specified number of iterations have been completed.
+After the trajectories are collected, the process proceeds to the policy update stage, where the trajectories stored in the replay buffer are used to calculate the total loss $L^{Total}$ for updating the policy and value networks.
+After updating the networks, the replay buffer will be cleared, and the agents collect trajectories again.
+The two stages alternate until convergence or episode limit is reached.
+Algorithm 1 shows a more detailed training process.
+In lines 3-9, each agent collects the trajectories respectively and calculates the values using the value network under the same environment.
+After specific steps, in lines 10-15, according to the trajectories collected by each agent, the corresponding advantages $A$ of each trajectory are calculated through General Advantage Estimation (GAE) [32].
+Then, we merge and shuffle the trajectories to get training data.
+In lines 17-24, we calculate the loss by:
+$\begin{array}[]{l}L^{{P}}=-\text{min}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)}A,\text{clip}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)},1-\epsilon,1+\epsilon)A),\\ L^{V}=\max((V_{\theta}-V_{targ})^{2},(\operatorname{clip}(V_{\theta},V_{\theta_{old}}-\varepsilon,\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad V_{\theta_{old}}+\varepsilon)-V_{tart})^{2}),\\ L^{{Entropy}}=H(\pi_{\theta}(a^{t}|o^{t})),\end{array}$ (1)
+where $V_{targ}=A+V_{\phi}(s)$.
+We get the total loss by weighted summing $L^{{Total}}=L^{{P}}+\alpha L^{V}+\beta L^{{Entopy}}$, where $\alpha$ and $\beta$ are hyperparameters.
+We use the Adam Optimizer [33] to update the policy and value network w.r.t $L^{Total}$.
+The algorithm proceeds to the next iteration until a specified iteration is reached.
+We also use some tricks to make the training more stable.
+The tricks include reward scaling [34], mini-batch updating, advantage normalizing, and learning rate decay.
+We use two scenarios during training: single-agent and multi-agent.
+As we clip the distance input to the network, the size of the scenario does not affect the training.
+Thus, in the single-agent scenario, the initial point and the goal point of navigation are 10 m apart, and the multi-agent scenario’s initial and goal points are 16 m apart.
+The two scenarios are shown in Fig.
+3.
+(a) Single-agent scenario
+(b) Multi-agent scenario
+[…]
+[…]
+The measurement dimension $L_{n}$ of the LiDAR of the policy is different, and the max length of the LiDAR beam is 4 m.
+We first train the policy with $L_{n}=36$, which works successfully in the single-agent scenario.
+In the multi-agent scenario, the $L_{n}$ is updated to 72.
+Because if $L_{n}=36$, the agent may lose information when other agents stay in the internal of the LiDAR beam, with the biggest internal = 0.34 m, while $L_{n}=72$, the biggest internal = 0.17 m is enough for most situations.
+The field of view (FOV) is set to $\pi$ rad.
+Additionally, we stack the LiDAR observations of 5 consecutive frames to provide more comprehensive observation information.
+The agent’s velocity is set within the $[0,1]$ m/s, and the angular velocity is set within the $[-\pi,\pi]$ rad/s.
+The control frequency used in training is 60 Hz, which can make the reward variation of each action more apparent and maintain efficient control.
+Finally, the maximum number of simulation steps is set to 2500, sufficient for the agent to finish navigation through the optimal path.
+III-D Reward Function
 The dense reward function used in RL-based navigation algorithms consists of distance reward, obstacle penalty, and efficiency reward. An example of such a reward function is:
-$\begin{array}[]{l}R=r_{g}+r_{c}+r_{other},\\ r_{g}=\left\{\begin{array}[]{ll}r_{{arrival}}\qquad\qquad\quad\ \ \text{if }\left\|p_{i}^{t}-g_{i}\right\|<0.1,\\ w_{g}\left(\left\|p_{i}^{t-1}-g_{i}\right\|-\left\|p_{i}^{t}-g_{i}\right\|\right)\quad\text{otherwise,}\end{array}\right.\\ r_{c}=\left\{\begin{array}[]{ll}r_{{collision}}\qquad\qquad\quad\text{if }d_{i}^{t}<0,\\ w_{c}(0.1-d_{i}^{t}/2)\quad\text{ if }d_{i}^{t}<0.2,\\ 0\qquad\qquad\qquad\quad\text{ otherwise,}\end{array}\right.\\ \end{array}$ (1)
-where $r_{other}$ is the other reward that varies in different approaches, $r_{g}$ means the distance reward, which is computed as the difference in the distance between the robot and the goal point at the current time step and the previous time step, $r_{c}$ means the obstacle penalty, $w_{g}$ and $w_{c}$ are the weight of $r_{g}$ and $r_{c}$ respectively, $d_{i}^{t}$ means the smallest distance between agent $i$ and obstacles at time step $t$.
+$\begin{array}[]{l}R=r_{g}+r_{c}+r_{other},\\ r_{g}=\left\{\begin{array}[]{ll}r_{{arrival}}\qquad\qquad\quad\ \ \text{if }\left\|p_{i}^{t}-g_{i}\right\|<0.1,\\ w_{g}\left(\left\|p_{i}^{t-1}-g_{i}\right\|-\left\|p_{i}^{t}-g_{i}\right\|\right)\quad\text{otherwise,}\end{array}\right.\\ r_{c}=\left\{\begin{array}[]{ll}r_{{collision}}\qquad\qquad\text{if }d_{i}^{t}<0.01,\\ w_{c}(0.1-d_{i}^{t})\qquad\text{ if }d_{i}^{t}<0.1,\\ 0\qquad\qquad\qquad\quad\text{ otherwise,}\end{array}\right.\\ \end{array}$ (2)
+where $r_{other}$ is the reward that varies in different approaches, $r_{g}$ is the distance reward, which is computed as the difference in the distance between the robot and the goal point at the current time step and the previous time step, $r_{c}$ is the obstacle penalty, $w_{g}$ and $w_{c}$ are the weight of $r_{g}$ and $r_{c}$ respectively, $d_{i}^{t}$ means the smallest distance between agent $i$ and obstacles at time step $t$.
 However, this reward function setup may make agents learn conservative policy.
@@
 (b) Multi-agent traveling in opposite directions.
-Figure 1: Cases that are difficult to handle with the classical reward function.
+Figure 4: Scenarios where the classical reward function demonstrates significant limitations.
+As shown in Fig.
+4, when an agent needs to navigate through the gap between two obstacles, it may incur penalties due to $r_{c}$ when approaching them, even though traversing the gap is the optimal action.
+As a result, the agent may circumnavigate obstacles.
+Furthermore, agents focus exclusively on proximal threats in multi-agent scenarios while ignoring potentially hazardous agents at greater distances, resulting in unsafe navigation decisions.
+To address the issues above, we have introduced modifications to the $r_{c}$ component, specifically by:
+$\begin{array}[]{l}r^{\prime}_{c}=\left\{\begin{array}[]{ll}r_{collision}&\text{if }d_{i}^{t}<0.01,\\ w^{\prime}_{c}\left(l-sl_{i}^{t}\right)+r_{c}&\text{ otherwise,}\end{array}\right.\end{array}$ (3)
+where $r_{arrival}=5$, $r_{collision}=-5$, $w_{g}=5$, $w^{\prime}_{c}=-0.25$, $w_{c}=-0.03$, $l$ represents the maximum length of the LiDAR rays, $sl_{i}^{t}$ represents the LiDAR observations of agent $i$ at time step $t$, $w^{\prime}_{c}$ is a weight vector that varies with angular velocity, with the weight decreasing in $\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-\omega\Delta t)^{2}}{2\sigma^{2}}}$ fashion from $\omega\Delta t$ towards both sides, where $\sigma=0.2$, $\Delta t$ is the time interval of the control.
+The schematic diagram of $w^{\prime}_{c}$ is shown in Fig.
+5.
 Figure 5: The weight of the scan penalty of the $r^{\prime}_{c}$.
-Figure 3: The overall framework diagram of our approaches is presented.
-The approaches can be divided into trajectory collection and policy update.
-During trajectory collection, each agent collects the trajectory independently and stores it in a replay buffer.
-The simulation platform will return each agent’s reward and observation after an action is executed.
-After a specific number of iterations, the trajectories stored in the replay buffer will be used to calculate the loss, and the network will then be updated based on the loss through gradient descent.
-The two stages alternate until a specific number of episodes is reached or the policy converges.
-As shown in Figure 1, when an agent needs to navigate through the gap between two obstacles, it may incur penalties due to $r_{c}$ when approaching them, even though traversing the gap is the optimal action.
-Consequently, the agent may choose to circumnavigate around the obstacles instead.
-Besides, when multiple agents approach each other, they may only focus on the closest agent or obstacle to itself, neglecting potentially threatening agents that are further away but pose a real danger, which leads to unsafe navigation.
-To address the issues above, we have introduced modifications to the $r_{c}$ component, specifically by:
-$\begin{array}[]{l}r^{\prime}_{c}=\left\{\begin{array}[]{ll}r_{collision}&\text{if }d_{i}^{t}<0,\\ w^{\prime}_{c}\left(l-sl_{i}^{t}\right)+r_{c}&\text{ otherwise,}\end{array}\right.\end{array}$ (2)
-where $r_{arrival}=5$, $r_{collision}=-5$, $w_{g}=5$, $w^{\prime}_{c}=-0.25$, $w_{c}=-0.03$, $l$ represents the maximum length of the LiDAR rays, $sl_{i}^{t}$ represents the LiDAR information of agent $i$ at time step $t$, $w^{\prime}_{c}$ is a weight vector that varies with angular velocity, with the weight decreasing in $\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-\omega\Delta t)^{2}}{2\sigma^{2}}}$ fashion from $\omega\Delta t$ towards both sides, where $\sigma=0.2$, $\Delta t$ is the time interval.
-The schematic diagram of $w^{\prime}_{c}$ is shown in Figure 2.
 Using $r^{\prime}_{c}$, the policy pays more attention to the direction of the current angular velocity while maintaining the closest distance to the surrounding obstacles.
-This allows the policy to better cope with the mentioned situations and allows the agent to pay more attention to the choice of angular velocity.
-4 Method
-In this section, we will introduce the architecture of the policy network, the simulation platform, and the training process with training scenarios.
-We utilize the hidden state of LSTM to compensate for the lack of information in the observations and propose an attention structure to improve the policy ability in the multi-agent scenario.
-Besides, we propose a multi-agent simulation platform to bridge the gap between simulation and the real world.
-Additionally, the multi-agent training process is based on Proximal Policy Optimization (PPO, Schulman et al.
-(2017)), with modifications in trajectory collection.
-Finally, we use two scenarios for training: single-agent and multi-agent.
-4.1 Network structure
-We note that the agent only requires more biased actions toward the goal point when approaching it.
-In other situations, the agent merely needs to move in a general direction.
-As such, we clip the distance between the agent and the goal point to enable the policy to better distinguish between these two situations rather than directly using the distance as an input to the network.
-Algorithm 1 PPO For Multi-Agent training
-1:procedure PPO(environment, hyperparameters)
-2:     Initialize policy network $\pi_{\theta}$, value function network $V_{\phi}$, and set hyper-parameters
-3:     for $iteration=1,2,...,$ do
-4:         for $t=1,2,...,T_{max}$ do
-5:              for robot $i=1,2,...$ do
-6:                  Run policy network $\pi_{\theta}$, collecting $\{o_{i}^{t},a_{i}^{t},r_{i}^{t}\}$.
-7:                  Run value function network, collecting $\{V_{\phi}(o_{i}^{t})\}$
-8:              end for
-9:         end for
-10:         for robot $i=1,2,...$ do
-11:              for $t=T_{max},T_{max-1},...,1$ do
-12:                  Calculate TD-error: $\delta_{i}^{t}=r_{i}^{t}+\gamma V_{\phi}(o_{i}^{t+1})-V_{\phi}(o_{i}^{t})$
-13:                  Estimate advantages using GAE: $A_{i}^{t}=\sum_{t}^{T_{max}}(\gamma\lambda)^{t}\delta_{i}^{t}$
-14:                  Calculate $V_{i}^{t}=A_{i}^{t}+V_{\phi}(o_{i}^{t})$
-15:              end for
-16:         end for
-17:         Merge and shuffle the collected trajectories
-18:         $\theta_{old}=\theta,\ \phi_{old}=\phi$
-19:         for epoch in Epoch do
-20:              $L^{{PPO}}=-\mathbb{E}\left(\sum_{i=1}^{N}\sum_{t=1}^{T_{max}}\text{min}\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{{old}}}(a_{i}^{t}|o_{i}^{t})}A_{i}^{t},\text{clip}\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{{old}}}(a_{i}^{t}|o_{i}^{t})},1-\epsilon,1+\epsilon\right)A_{i}^{t}\right)\right)$
-21:              $L^{V}=\text{max}\left(\mathbb{E}\left(\sum_{i=1}^{N}\sum_{t=1}^{T_{max}}\left((V_{\phi}(o^{t}_{i})-V_{i}^{t})^{2},\left(\text{clip}\left(V_{\phi}(o_{i}^{t}),V_{\phi_{old}}(o_{i}^{t})-\epsilon,V_{\phi_{old}}(o_{i}^{t})+\epsilon\right)-V_{i}^{t}\right)^{2}\right)\right)\right)$
-22:              $L^{{Entropy}}=H(\pi_{\theta}(a^{t}|o^{t}))$
-23:              $L^{{Total}}=L^{{PPO}}+\alpha L^{V}+\beta L^{{Entopy}}$
-24:              Update the parameters of $\pi_{\theta}$ and $V_{\phi}$ using Adam Optimizer w.r.t $L^{Total}$
-25:         end for
-26:     end for
-27:end procedure
-We first only use LSTM to encode the LiDAR observation.
-The policy is trained in a static environment and works well in static and 6-agent test environments.
-However, it performs poorly when the robot’s number increases over 6.
-Thus, we add an attention module to overcome the issue.
-As shown in Figure 3, we use a three-layer LSTM module to encode the laser measurements $o_{s}^{t}$, with both hidden sizes = 256, and get $S$.
-Then, $K$ and $V$ are calculated by two different FC layers by using $S$, in which the output sizes = 256.
+The reward based on angular velocity of the agent allows the policy to better cope with the mentioned situations and enable the agent to pay more attention to the choice of angular velocity.
+III-E Network Structure
+The agent only requires more precise, goal-oriented actions when nearing the target location, whereas general directional movement suffices otherwise.
+We clip the agent-goal distance rather than feeding raw distance values into the network to enable the policy to differentiate between these two regimes explicitly.
+This design decouples the policy’s decision-making from absolute distance metrics, ensuring scalability to larger environments while maintaining consistent behavioral logic across spatial scales.
+To address the loss of information in POMDP, we first employ LSTM Layers to encode LiDAR observations.
+By stacking sequential LiDAR observations, the LSTM module captures historical observation memory and integrates consecutive observation patterns.
+While the policy performs well in static environments and 6-agent test scenarios, its performance degrades significantly when the number of agents exceeds 6.
+This limitation arises because the LSTM-based policy overly focuses on LiDAR observations while neglecting the agent’s velocity, making it difficult to navigate effectively in denser multi-agent environments.
+To address this issue, we introduce an attention mechanism that dynamically balances the importance of LiDAR observations and velocity information.
+First, the velocity and goal information are encoded through fully connected (FC) layers to produce the query vector $Q$, while the LiDAR observations are processed by the LSTM module with FC layers to generate the value $V$ and key $K$ vectors, respectively.
+The attention mechanism then integrates all relevant information into a comprehensive representation, as described in (4).
+Integrating the proposed reward function enhances policy and achieves robust navigation in complex multi-agent scenarios.
+The architecture, which combines LSTM with an attention mechanism, is designated the LSTM-Attention policy.
+TABLE I: Experiment with 1 robot and 15 obstacles.
+success rate(%) collision rate(%) trap rate (%) average step
+IJRR [12] 91.3 8.6 0.1 1383.23
+  Linear policy 67.2 2.4 30.4 550.40
+LSTM policy 93.1 6.9 0.0 1187.73
+LSTM-Attention 95.9 3.9 0.2 1300.78
+  (Safe) LSTM-Attention 97.6 2 0.04 1353.63
+TABLE II: Experiment with 6 robots and 5 obstacles.
+success rate(%) collision rate(%) trap rate (%) average step
+NH-ORCA [6] 98.4 0.33 1.27 1198.82
+IJRR [12] 85.03 14.63 0.33 1617.84
+LSTM policy 91.89 8.15 0.0 981.01
+LSTM-Attention 94.78 4.68 0.53 1805.94
+As shown in Fig.
+2, we use a three-layer LSTM module to encode the laser measurements $o_{s}^{t}$ with both hidden sizes = 256 and get LSTM encoding $S$.
+Then, $K$ and $V$ are computed using two FC layers, each with the output sizes = 256, applied to the encoding $S$.
 At the same time, $[o_{g}^{t},\ o_{v}^{t}]$ passes through two FC layers, which output sizes = 256, to obtain $Q$.
-Then, the encoding of $o_{s}^{t}$, $X$, is obtained by multi-head attention:
+Then, the encoding of the LiDAR observations, $X$, is obtained by multi-head attention:
 $\begin{array}[]{l}\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h})W^{O},\\ \text{where }\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),\\ \text{Attention}(Q,K,V)=\text{Softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V,\end{array}$ (4)
 where $d_{k}$ is the size of $K$.
-The final encoding is obtained by the weighted sum of $X$ and $S$ with an adaptive residual adjustment, $\alpha X+(1-\alpha)S$, where $\alpha$ updates with the policy and the value network.
-In actor and critic, the final encoding is decoded by two fully connected layer (FC) layers with output size = 128 and output size = 2 to get the mean of the policy $[v^{t},\omega^{t}]$, and another two FC layers with output size = 128 and output size = 1 to get the value network $V_{\phi}(o^{t})$.
+The final encoding is obtained by the weighted sum of $X$ and $S$ with an adaptive residual adjustment, $\alpha X+(1-\alpha)S$, where $\alpha$ updates with the network.
+In actor and critic, the final encoding is decoded by two FC layers with output size = 128 and output size = 2 to get the mean of the policy $[v^{t},\omega^{t}]$, and another two FC layers with output size = 128 and output size = 1 to get the value network $V_{\phi}(o^{t})$.
 Note that each LSTM layer is followed by a rectified linear unit (ReLU), and so is the FC layer.
 Additionally, the last FC layer of the actor and critic modules has no ReLU layers.
-4.2 Simulation platform
-To bridge the gap between simulation and the real world, we propose a multi-agent training platform based on a physics-based simulator such as Pybullet (Coumans and Bai (2016–2021)) and Gazebo.
-In most 2-D simulators, the actions of the agent are usually perfectly executed, and their positions are usually calculated by $x^{t+1}=x^{t}+v^{t}\Delta t+\frac{1}{2}a^{t}\Delta t^{2}$.
-In a physics-based simulator, the agent state is computed by the physics engine, i.e., $x^{t+1}=T(x^{t},v^{t},a^{t},f^{t},...)$, where $T$ is the transfer function of the physics engine, $f^{t}$ is the force acting on the agent.
-The transfer function $T$ makes the execution of the action randomness.
-A straightforward example is when the straight action executes in different types of simulators, the results differ.
-As shown in Figure 4.
-The 2-D simulator executes the action without bias, while the agent in the physics-based simulator deviates from the straight line, and the deviation is greater in real-world robots.
-This indicates that training on a physics-based simulator may help the agent learn more realistic policies, thereby facilitating the deployment of the policy on actual robots.
-To better train the policy, we add Gaussian noise to the agent’s perception to increase randomness further.
-(a) 2-D simulator.
-(b) 3-D simulator.
-(c) Real robot.
-Figure 4: Comparison between the simulators and the real world.
-Agents on all platforms execute commands with $v=1\ m/s,\ \omega=0\ rad/s$.
-The direct travel distance of the agent in the simulator is 10m, and the distance of the real robot is 5 m.
-We propose a local replay mechanism for the trajectory collection of the agent’s collision aiming to speed up the collection of multi-agent interaction trajectories.
-When the agent collides, the agent will be reset back to the state before the $N$ simulation steps so that the agent can quickly accumulate the collision data and speed up the training, where $N=300$ in this article.
-However, when the local replay mechanism is deployed in a multi-agent scenario, the reset agent may appear in front of the normal agent, causing the normal agent to be disturbed.
-Sudden disturbances may make the policy conservative.
-However, we found that in some cases, agents perform slowing or stopping actions to wait for other agents to leave.
-Besides, to make the simulator generalize, we adopt the architecture of Gym (Brockman et al.
-(2016)), and we have diligently maintained the interfaces intending to provide users with a more convenient experience when utilizing the simulator.
-The simulator supports the random addition of random or queued agents or obstacles and controls the movement of each agent through a action set $\mathcal{A}$.
-At the end of a simulator step, the simulator returns a observed state set $\mathcal{O}$.
-The update frequency of the simulator can be modified to meet the policy training at different control frequencies.
-All of these functionalities are designed for easy reconfiguration.
-For example, to accommodate users seeking to evaluate Safe Reinforcement Learning (Safe-RL, Achiam and Amodei (2019)), we have reconfigured our custom-designed simulator to include a constrained version by changing two functions of the origin version.
-4.3 Training process
-We treat distributed multi-agent navigation as plenty of single agents navigating in an unknown environment, meaning each agent collects the trajectory respectively.
-As shown in Figure 3, the training process can be divided into trajectory collection and policy update.
-During the trajectory collection, each agent first observes the LiDAR $o_{s}^{t}$, target point $o_{g}^{t}$, and velocity $o_{v}^{t}$.
-After that, it utilizes the policy and value network to calculate a value $v^{t}$ and an action mean $[v^{t},\omega^{t}]$.
-From these, an action $a^{t}$ is sampled based on the policy distribution $\pi_{\theta}(a^{t}|o^{t})$ and executed.
-The simulator then responds with a reward $r^{t}$, the next state $o^{t+1}$, and a termination signal $done^{t}$.
-The tuple of $<o^{t},a^{t},r^{t},o^{t+1},done^{t},V_{\phi}(o^{t})>$ is stored in replay buffer.
-This process is repeated until a specified number of iterations have been completed.
-Subsequently, the trajectories stored in the relay buffer are used to calculate the total loss $L^{Total}$ for updating the policy and value networks.
-After updating the networks, the replay buffer will be cleared, and the agents collect trajectories again.
-The two stages alternate until convergence or episode limit is reached.
-Alg 1 shows a more detailed training process.
-In lines 3-9, each agent collects the trajectories respectively and calculates the values using the value network under the same environment.
-After specific steps, in lines 10-15, according to the trajectories collected by each agent, the corresponding advantages of each trajectory are calculated through General Advantage Estimation (GAE, Schulman et al.
-(2015)).
-Then, we merge and shuffle the trajectories to get training data.
-In lines 17-24, we calculate the policy loss by $L^{{PPO}}=\text{min}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)}A,\text{clip}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)},1-\epsilon,1+\epsilon)A)$, the value loss by $L^{V}=\max((V_{\theta}-V_{targ})^{2},(\operatorname{clip}(V_{\theta},V_{\theta_{old}}-\varepsilon,V_{\theta_{old}}+\varepsilon)-V_{tart})^{2})$, where $V_{targ}=A+V_{\phi}(s)$ and the entropy loss by $L^{{Entropy}}=H(\pi_{\theta}(a^{t}|o^{t}))$.
-We get the total loss by weighted summing $L^{{Total}}=-L^{{PPO}}+\alpha L^{V}+\beta L^{{Entopy}}$, where $\alpha$ and $\beta$ are hyperparameters.
-We use the Adam Optimizer (Kingma and Ba (2014)) to update the policy and value network w.r.t $L^{Total}$.
-The algorithm proceeds to the next iteration until a specified iteration is reached.
-We also use some tricks to make the training more stable.
-The tricks include reward scaling (Gao et al.
-(2022)), mini-batch updating, advantage normalizing, and learning rate decay.
-We use two scenarios during training: single-agent and multi-agent.
-As we clip the distance input to the network, the size of the scenario does not affect the training.
-Thus, in the single-agent scenario, the initial point and the goal point of navigation are 10 m apart and the multi-agent scenario’s initial and goal navigation points are 16 m apart.
-The two scenarios are shown in Figure 5.
-(a) Single-agent scenario.
-(b) Multi-agent scenario.
-[…]
-[…]
-Table 1: Experiment with 1 robot and 15 obstacles.
-success rate(%) collision rate(%) trap rate (%) average step
-IJRR (Fan et al. (2020)) 91.3 8.6 0.1 1383.23
-  Linear policy 67.2 2.4 30.4 550.40
-LSTM policy 93.1 6.9 0.0 1187.73
-LSTM+Attention 95.9 3.9 0.2 1300.78
-  (Safe) LSTM+Attention 97.6 2 0.04 1353.63
-Table 2: Experiment with 6 robots and 5 obstacles.
-success rate(%) collision rate(%) trap rate (%) average step
-NH-ORCA (Alonso-Mora et al. (2010)) 98.4 0.33 1.27 1198.82
-IJRR (Fan et al. (2020)) 85.03 14.63 0.33 1617.84
-LSTM policy 91.89 8.15 0.0 981.01
-LSTM+Attention 94.78 4.68 0.53 1805.94
-Figure 6: Multi-agent test without obstacle.
-The measurement dimension $L_{n}$ of the LiDAR of the policy is different, and the max length of the LiDAR beam is 4 m.
-We first train the policy with $L_{n}=36$, which works successfully in the single-agent scenario.
-In the multi-agent scenario, the $L_{n}$ of LSTM policy and LSTM+Attention are 72.
-Because, if $L_{n}=36$, the agent may lose information when other agents stay in the internal of the LiDAR beam, with the biggest internal = 0.34 m, while $L_{n}=72$, the biggest internal = 0.17 m is enough for most situations.
-The field of view (FOV) is set to $\pi$ rad.
-Additionally, we stack the LiDAR observations of 5 consecutive frames to provide more comprehensive observation information.
-The agent’s velocity is set within the $[0,1]$ m/s, and the angular velocity is set within the $[-\pi,\pi]$ rad/s.
-The control frequency used in training is 60 Hz, which can make the reward variation of each action more apparent and maintain efficient control.
-Finally, the maximum number of simulation steps is set to 2500, sufficient for the agent to finish navigation through the optimal path.
-5 Experiments
-To evaluate the ability of the policy, we test it in both simulation and the real world compared with NH-ORCA (Alonso-Mora et al.
-(2010)) and the policy of Fan et al.
-(2020).
-In the following section, we refer to the policy of Fan et al.
-(2020) by CNN-policy, and the measurement dimension of the LiDAR $L_{n}=512$.
-The LSTM+Attention policy is the proposed policy.
-We conducted an ablation experiment by removing the attention mechanism from the proposed policy, which is LSTM policy.
-Additionally, we replaced the LSTM layer with a Linear layer and get the Linear policy, in order to validate the effectiveness of the LSTM layer.
-In the simulation tests, we use some metrics to evaluate the policy.
+IV EXPERIMENT EVALUATION
+IV-A Experimental Setup
+All models are implemented with PYTORCH on a server with a GTX4090 GPU and CPU of Intel Xeon 8336C. We test the model and select their optimal parameter settings.
+To evaluate the ability of the policy, we test it in both simulation and the real world compared with NH-ORCA [6] and the policy of [12].
+In the following section, we refer to the policy of [12] by the CNN policy and the measurement dimension of the LiDAR $L_{n}=512$.
+The LSTM-Attention policy is the proposed policy.
+We conduct an ablation experiment by removing the attention mechanism from the proposed policy, which is the LSTM policy.
+Additionally, we replace the LSTM layer with the Linear layer and obtain the Linear policy to validate the effectiveness of the LSTM layer.
+In the simulation tests, we use four metrics to evaluate the policy.
 The metrics consist of the success rate, the collision rate, the trap rate, and the average step.
@@
 The average step is the average simulation step of the agent to reach the goal successfully.
-We use the simulation step rather than the cost time of navigation because the cost time depends on the calculation speed of the simulator, which cannot exactly reflect the cost time of navigation.
-Finally, all simulation tests are performed 1000 times.
-5.1 Simulation test
+We use simulation steps instead of navigation cost time as the metric because the latter depends on the computational speed of the simulator, which may not accurately reflect real-world navigation time.
+Finally, we perform all simulation tests 1000 times.
+Figure 6: Multi-agent test without obstacle.
+IV-B Simulation Test
 […]
 […]
+Figure 7: Real-world single-agent experiment.
+The TurtleBot3 first passes through narrow passages, crosses through the obstacles, and finally reaches the goal point.
+The TurtleBot3 reaches the goal point in 46 s.
+IV-C Real-World Test
+To evaluate the policy’s performance in real-world scenarios, we deploy it on a TurtleBot3 platform.
+The TurtleBot3’s maximum linear and angular velocities are limited to 0.2 m/s and 1 rad/s, respectively; thus, we scale the policy’s output to align with these constraints.
+Additionally, the control frequency is set to 10 Hz.
+The LiDAR sensor on the TrutleBot3 provides FOV $=2\pi$ rad with 360 beams, but we only utilize 72 beams within a reduced FOV $=0.8\pi$ rad to match our experimental requirements.
+We align the rest of the policy with the simulation settings.
+We use the wheel odometry and LiDAR sensor on the TurtleBot3 for perception.
+Wheel odometry calculates the robot’s relative position and orientation to the goal, while the LiDAR sensor captures the surrounding environment observations.
+In the single-agent scenario, we evaluate the policy’s performance by navigating the TurtleBot3 through narrow passages.
+The TurtleBot3 successfully passes the passages and avoids obstacles of unknown positions and shapes while reaching the goal point.
+As illustrated in Fig.
+7, the TurtleBot3 begins in a room, traverses a narrow passage to enter the goal room in 22 s, and then avoids obstacles to achieve the goal point in an additional 24 s.
+Finally, the TurtleBot3 completes the navigation task as it approaches the goal point.
+In the multi-agent scenario, we evaluate the policy’s performance in avoiding other agents and responding to sudden disturbances.
+Each TurtleBot3 needs to navigate through obstacles while avoiding the other robots and human.
+In this evaluation, the human acts as a dynamic disturbance by intentionally interfering with the TurtleBot3s.
+As shown in Fig.
+8, the human first appears in front of a TurtleBot3 to force it to alter its trajectory, then crosses directly between two TurtleBot3s.
+Subsequently, the human continues interfering with one TurtleBot3 until it reaches its goal point.
+Despite these disturbances, all TurtleBot3s successfully avoid collisions with each other and the human.
+The navigation algorithm demonstrates robust performance in both single-agent and multi-agent scenarios.
+In the single-agent scenario, the TurtleBot3 successfully navigates through narrow passages and avoids obstacles of unknown positions and shapes, efficiently reaching the goal point.
+In multi-agent scenario, the algorithm enables TurtleBot3s to avoid collisions with other robots and respond to dynamic disturbances, such as intentional human interference.
+The TurtleBot3s dynamically adjust their trajectories to accommodate sudden obstacles while maintaining collision-free navigation.
+These results highlight the algorithm’s adaptability and reliability in handling static and dynamic challenges, ensuring safe and efficient navigation in real-world applications.
 Figure 8: Real-world multi-agent experiment.
@@
 All TurtleBot3s reach the goal points in 31 s.
-Figure 8: Real-world single-agent experiment.
-The turtlebot3 first passes through narrow passages, then crosses through the obstacles, and finally reaches the goal point.
-The turtlebot3 reaches the goal point in 46 s.
-5.2 Real-world test
-We use turtlebot3 to test the policy’s performance in the real world.
-As the max linear velocity and angular velocity of the turtlebot3 are 0.2 m/s and 1 rad/s, we scale the output of our policy to fit the ability of turtlebot3.
-Besides, the control frequency of turtlebot3 is set to 10 Hz.
-The observation of the LiDAR given by turtlebot3 with parameters are FOV $=2\pi$ rad, and LiDAR beam number = 360.
-We only use 72 LiDAR beam with FOV $=0.8\pi$ rad.
-We align the rest of the policy with the simulation settings.
-We use the wheel odometry and LiDAR sensor on the turtlebot3 for perception.
-Wheel odometry calculates the robot’s relative position and angle to the goal point, while the LiDAR sensor is employed to gather LiDAR observation.
-Although it is possible to let each turtlebot3 calculate its action independently, it is unable to install the corresponding library, like Pytorch Paszke et al.
-(2017), on the onboard Raspberry Pi.
-Therefore, we control turtlebot3 through the Robot Operating System (ROS, Quigley et al.
-(2009)) communication protocol by using a host computer.
-In the single-agent scenario, we test the performance of the policy through narrow passages.
-The turtlebot3 passes and avoids obstacles with unknown positions and shapes to reach the goal point.
-As shown in Figure 8, the turtlebot3 starts at a room and passes through a passage to get in the room with the goal point in 22 s.
-Then, the turtlebot3 crosses the obstacles to reach the goal point in 24 s.
-At last, the turtlebot3 finishes the navigation when approaching the goal point.
-In the multi-agent scenario, we test the policy’s performance in avoiding other agents and sudden disturbances.
-A turtlebot3 needs to go through obstacles and avoid other turtlebot3.
-Besides, a human who wants to interfere with turtlebot3 is a sudden disturbance.
-As shown in Figure 7, at first, the human appears in front of a turtlebot3 to force it to change its trajectory, and after that, the human directly crosses through two turtlebot3.
-Then, the human continues to interfere with other turtlebot3 until it reaches its goal point.
-All turtlebot3s successfully avoid collision with other turtlebot3s and the human.
-These two experiments prove that the policies trained on our proposed simulation platform can be directly deployed on real robots, with only adjustments made to the values of output actions based on the real robot’s capabilities.
-6 Conclusion
-This paper proposes a lightweight, end-to-end multi-agent navigation policy that translates raw LiDAR observation directly into action commands, eliminating the need to distinguish between dynamic and static obstacles and enabling navigation through narrow gaps.
-A physics-engine-based simulation platform is introduced for multi-agent training, bridging the gap between simulation and reality.
-Our method, trained using RL, demonstrates robust performance in simulated and real-world environments, even in malicious obstructions.
-Key contributions include the proposal of a distributed navigation policy trained in stochastic environments, the utilization of LSTM to mitigate information loss in POMDP, the development of a multi-agent training platform to facilitate idea validation and safe RL, and the introduction of a new reward function to address suboptimal action selection.
-The policy’s performance is evaluated in simulation and real-world settings, highlighting its effectiveness and potential for future enhancements.
-In the future, we will explore extending the policy to more complex environments, incorporating passing through traffic, or expanding the work to Unmanned aerial vehicles (UAVs).
-We will further refine the reward function and the policy to enhance performance in edge cases, like U-shaped obstacles.
-{acks}
-This work is supported by Shenzhen Science and Technology Program under Grant RCBS20221008093305007, Grant 20231115141459001, Young Elite Scientists Sponsorship Program by CAST under Grant 2023QNRC001, and High level of special funds (G03034K003) from Southern University of Science and Technology, Shenzhen, China.
+V CONCLUSION AND FUTURE WORK
+We present a lightweight, end-to-end multi-agent navigation framework that directly maps raw LiDAR observations to action commands, eliminating the need for explicit obstacle classification while enabling efficient navigation through constrained spaces.
+Our approach leverages a physics-enhanced simulation platform to bridge the simulation-to-real gap, facilitating practical multi-agent training.
+The reinforcement learning-based solution demonstrates robust performance across simulated and real-world environments, including challenging scenarios with adversarial obstructions, while maintaining computational efficiency.
+Our work makes several key contributions to the field of multi-agent navigation.
+We develop a distributed navigation policy capable of operating in stochastic environments, incorporating an LSTM-based architecture to address information loss in POMDPs.
+The proposed physics-integrated training platform enables safe and efficient RL validation, complemented by a novel reward function that effectively resolves suboptimal action selection in complex scenarios.
+Comprehensive evaluations in diverse environments demonstrate the framework’s effectiveness, robustness, and real-world applicability.
+In the future, we will identify multiple promising directions for future research.
+These include extending the framework to handle dynamic traffic scenarios and unmanned aerial vehicle (UAV) applications, refining the reward function to address complex edge cases such as U-shaped obstacles, and enhancing policy robustness in adversarial environments.
+The proposed framework’s adaptability and performance suggest strong potential for broader applications in autonomous navigation systems.